### v1

# Design an Editable Speech-to-Sign-Language Transformer System: A Human-Centered AI Approach

Abstract
This paper presents a human-centered, real-time, user-adaptive speech-to-sign language animation system that integrates Transformer-based motion generation with a transparent, user-editable JSON intermediate layer. The framework overcomes key limitations in prior sign language technologies by enabling direct user inspection and modification of sign segments, thus enhancing naturalness, expressiveness, and user agency. Leveraging a streaming Conformer encoder and autoregressive Transformer-MDN decoder, the system synchronizes spoken input into upper-body and facial motion for 3D avatar rendering. Edits and user ratings feed into a human-in-the-loop optimization loop for continuous improvement. Experiments with 20 deaf signers and 5 interpreters show that the editable interface and participatory feedback significantly improve comprehension, naturalness, usability, and trust, while lowering cognitive load. With sub-20 ms per-frame inference on standard hardware, the system is ready for real-time communication and education. This work illustrates how technical and participatory innovation together enable accessible, explainable, and user-adaptive AI for sign language technology.
sign language animation; human-centered AI; transformer; user-editable interface
1 Introduction
Approximately 70 million deaf individuals worldwide use sign language as their first language (WHO, 2021), yet they continue to face communication barriers in education, healthcare, and public services. Most mainstream sign language assistance systems rely on a multi-stage pipeline of speech → text → gloss → pre-recorded animations, often resulting in rigid outputs lacking facial and torso signals. Additionally, users struggle to effectively modify or personalize AI-generated results, leading to low adoption rates within the deaf community (Dimou et al., 2022).
In recent years, the Transformer architecture has emerged as a core method in motion generation and sign language synthesis due to its exceptional performance in sequence modeling (Saunders et al., 2020b). Transformers can effectively capture long-range dependencies, enabling direct mapping from speech or text to continuous 3D keypoint sequences. However, existing approaches still face two critical challenges: first, models often function as ”black boxes,” lacking interpretability and user engagement mechanisms; second, constrained by computational complexity and pipeline coupling, practical systems frequently struggle to achieve low-latency, highly interactive deployment (Saunders et al., 2020a).
To address these challenges, this paper adopts a Human-Centered AI (HCAI) framework, proposing an end-to-end streaming Conformer-Transformer pipeline that directly converts speech or text into 3D sign language motions (including upper-body and facial keypoints). The system enhances usability, interpretability, and user trust through structured, editable JSON scripts and a human-in-the-loop feedback mechanism. Our design empowers users—particularly deaf individuals and sign language interpreters—to intuitively review, edit, and personalize each sign language sequence via drag-and-drop interfaces and real-time animation playback, achieving truly ”controllable” AI output. The system continuously collects user edits and ratings, periodically fine-tuning models with this feedback to ensure alignment with real-world needs.
To systematically validate the effectiveness of human-centered design, we developed a 34-item Likert-scale questionnaire covering seven dimensions (comprehensibility, naturalness, interpretability, controllability, trust, etc.) and conducted comparative evaluations with 20 native deaf signers and 5 professional interpreters across two modes: ”Auto” (automatic generation) and ”Edit” (generation + editing). Results show that Edit mode improved comprehension by 27%, naturalness by 23%, and System Usability Scale (SUS) scores by 13 points, while reducing overall cognitive load by 15% (see Section 4).
The contributions of this paper are as follows:
  • Real-time streaming skeletal motion generation: Proposes an end-to-end sign language motion generation system based on streaming Conformer-Transformer, balancing high fidelity and low latency to meet practical accessibility needs (Saunders et al., 2020a; Damdoo & Kumar, 2025).
  • Structured editable interaction layer: Designs JSON scripts and a drag-and-drop editor covering elements such as gloss, handshape, duration, and facial expressions to enhance interpretability and controllability.
  • Human-centered optimization and closed-loop iteration: Introduces a continuous optimization mechanism based on user feedback and expert annotations, supporting model fine-tuning and adaptive system evolution.
  • Empirical evaluation: Through mixed quantitative and qualitative experiments, this study is the first to empirically validate the significant improvements in comprehension, naturalness, usability, and other multidimensional metrics through human-centered design.
The paper is structured as follows: Section 2 reviews related work; Section 3 describes the system architecture and data flow; Section 4 presents evaluation methods and results; Section 5 discusses limitations and improvements; Section 6 concludes with future work.
2 Relative Works
2.1 Transformer in Sign Language Motion Generation
The Transformer’s self-attention mechanism effectively captures long-range sequence dependencies. Since its introduction by Vaswani et al. (2017), it has been widely applied to motion generation and sign language modeling. Saunders et al.’s (2020b) Progressive Transformer treats sign language generation as a sequence-to-sequence translation task, directly mapping text or gloss sequences to continuous 3D skeletal point sequences, achieving state-of-the-art (SOTA) performance. Subsequent research improved naturalness and diversity through multi-channel modeling, Mixture Density Networks (MDN), and multi-task optimization (Saunders et al., 2020a). Damdoo and Kumar’s (2025) SignEdgeLVM achieved near 30 FPS real-time skeletal animation generation on edge devices with low computational power, demonstrating the lightweight potential of Transformer architectures.
Recent advancements have also been made in diffusion and latent variable Transformer methods: wSignGen achieved more realistic motion details and grammatical accuracy in word-level 3D ASL motion generation (Dong et al., 2024b); SignAvatar combined CVAE with a Transformer architecture to achieve highly robust 3D motion reconstruction and generation (Dong et al., 2024a); the latent variable Transformer proposed by Xie et al. (2024) surpassed traditional Seq2Seq models on both WLASL and PHOENIX14T benchmarks; SinDiff significantly improved the coherence and detail consistency of long-sequence sign language generation through a Transformer diffusion framework (Liang & Xu, 2023).
2.2 Sign Language Translation Paradigms
Traditional sign language translation systems often adopt a ”text→gloss→motion” intermediary approach, which is logically clear but requires extensive gloss annotation and struggles to capture facial and torso grammar (Tan et al., 2024). For convenient cross-comparison, commonly used benchmark datasets include RWTH-PHOENIX-Weather 2014T (Koller et al., 2015), WLASL (Li et al., 2020), and the phoneme-annotated WLASL-LEX (Tavella et al., 2022), which provide unified standards for evaluating the performance of different paradigm approaches. Gloss-free end-to-end methods reduce annotation needs through weak supervision or latent variable alignment: for example, GASLT (Yin et al., 2023) and GloFE (Lin et al., 2023) proposed weak supervision mechanisms based on gloss-attention and semantic alignment, respectively. SignVQNet (Hwang et al., 2024) used discretized latent codebooks to enable direct text-to-motion mapping but still faces challenges in data scale and temporal synchronization.
2.3 System Optimization
In practical deployment, generated sign language motions must drive avatars or animated characters in real time. Cui et al. (2022) proposed a 3D skeletal point regression method based on spatio-temporal graph convolution, combined with inverse kinematics (IK) for smooth animation. Shi et al. (2024) achieved significant improvements in inter-frame consistency and motion smoothness through their fine-grained video generation technology based on optical flow warping and pose fusion modules. Gan et al. (Gan et al., 2023) achieved end-to-end inference for 9.2s videos on edge devices, highlighting the importance of lightweight models and efficient rendering pipelines.
2.4 Human-Centered Design
Human-Centered AI (HCAI) emphasizes three principles: ”transparency, controllability, and trustworthiness” (Shneiderman, 2022), advocating for deep involvement of target users in AI system design, testing, and iterative feedback. The foundational work in interactive machine learning also supports HCAI theory: Fails & Olsen (2003) proposed the concept of ”Interactive Machine Learning,” emphasizing users’ active role in model training; Amershi et al. (2014) systematically summarized the crucial role of human-computer collaboration in interactive machine learning, further highlighting the importance of participatory feedback for improving model performance. Dimou et al. (2022) demonstrated that meaningful participation by Deaf communities significantly improves the acceptance of sign language avatars. Kothadiya et al.’s (2023) SignExplainer framework integrated explanation layers and user correction in recognition tasks, showing that ”explainable” design enhances AI trustworthiness. However, existing sign language generation systems rarely support real-time editing or human-AI collaborative closed-loop optimization. Our work fills this gap by systematically validating the multidimensional benefits of human-centered design principles in accessible sign language animation.
3 Methodology
Modern AI-powered sign language generation systems must not only achieve real-time efficiency and natural movements, but also deeply integrate human-centered design principles to fundamentally address users’ diverse needs and societal ethical expectations. This section details the overall architecture, algorithmic principles, and human-centered interaction flow of the end-to-end speech-driven sign language animation system proposed in this study. Our solution is designed around three core principles - ”real-time performance, explainability, and user participation” - achieving for the first time an organic integration of Transformer-generated motion sequences, structured intermediate representations, and user-controllable closed-loop optimization.
3.1 System Architecture and Data Flow
Our end-to-end speech-to-sign animation pipeline comprises six tightly integrated modules:
  1. 1.
    Streaming Conformer Encoder: Incoming audio frames $x_{1:T}$ are converted into high-level representations $\mathbf{H}$ with an internal encoder delay of $\leq 50ms$; together with decoding, IK, and rendering stages (see §3.4), the total *speech-to-avatar* latency is bounded to $\leq 150ms$ end-to-end.
  2. 2.
    Autoregressive Transformer-MDN Decoder: Conditions on $\mathbf{H}$ and previous motion latents $\{z_{<t}\}$ to produce a sequence of 128-dim latent vectors $z_{t}$, gloss labels $g_{t}$, and AU labels $a_{t}$.
  3. 3.
    Structured JSON Generator: Maps $\{z_{t},g_{t},a_{t}\}$ into a human-readable intermediate representation $\mathcal{J}$, exposing fields $\{\texttt{gloss},\texttt{start},\dots\}$.
  4. 4.
    Interactive JSON Editor: Allows users to inspect and modify $\mathcal{J}$; any edit triggers local resampling of the affected $z_{t}$ subsequence.
  5. 5.
    Unity3D IK Renderer: Binds the final motion latents $\{z^{\prime}_{t}\}$ to a 3D humanoid rig using Two-Bone IK and spline smoothing, producing real-time animation $\mathcal{A}$.
  6. 6.
    Edge-side Optimization & HITL Feedback: Applies model pruning and quantization for sub-20 ms/frame inference, while capturing user edits and ratings for periodic human-in-the-loop fine-tuning.
Figure 1: Detailed system architecture for our human‐centered sign language generation pipeline.
During initial design phases, three rounds of interviews and co-creation workshops with deaf users and interpreters surfaced three principal requirements:
  • Real-time alignment: end-to-end latency $<128$ ms,
  • Expressive diversity: synchronized upper-body and facial motion generation,
  • Full user agency: transparent, editable intermediate layer with continuous user intervention.
Following the IDEO “Insight–Principle–Solution” framework (IDEO.org, 2015), these requirements directly informed our three architectural pillars:
[…]
(iii) Editable JSON layer + human-in-the-loop optimization.
3.2 Streaming Conformer Encoder with Transformer-MDN Action Generation
The system first splits the input audio into frames (252525\,25ms window, 101010\,10ms hop) and extracts an $80$-dimensional Mel-spectrogram sequence
$X=\{x_{t}\}_{t=1}^{T}.$
A 6-layer streaming Conformer (each layer with model dimension $d=256$, 4-head causal self-attention plus local conv) (Gulati et al., 2020) then encodes $X$ into a downsampled prosody–semantic feature sequence
$H=\{h_{n}\}_{n=1}^{N},\quad N\approx T/r,\;r\in\mathbb{N},$
ensuring end-to-end latency $<160\,$ms (Damdoo & Kumar, 2025). Details are shown in Algorithm 1.
Algorithm 1 Streaming Conformer Encoder with State Caching
  Input: audio frame $x_{t}$, previous state $S_{t-1}$
  Output: feature $h_{t}$, updated state $S_{t}$
  // Extract Mel frame
  $f_{t}\leftarrow\mathrm{MelSpec}(x_{t})$
  // Linear projection
  $u_{t}\leftarrow W_{p}f_{t}+b_{p}$
  // Causal Conformer layers
  for $l=1$ to $L$ do
     $c_{t}^{(l)}\leftarrow\mathrm{ConvModule}^{(l)}(u_{t},S_{t-1}^{(l)})$
     $a_{t}^{(l)}\leftarrow\mathrm{CausalSelfAttn}^{(l)}(u_{t},S_{t-1}^{(l)})$
     $u_{t}\leftarrow\mathrm{LayerNorm}(u_{t}+c_{t}^{(l)}+a_{t}^{(l)})$
     Update $S_{t}^{(l)}$ with current keys/values
  end for
  $h_{t}\leftarrow u_{t}$
In the decoding stage, an autoregressive Transformer-MDN generates, at each step $t$ the autoregressive Transformer-MDN outputs a 128-dimensional latent vector $z_{t}\in\mathbb{R}^{128}$. This vector is produced by a two-stage VAE compressor that projects the 228 raw SMPL-X pose parameters (75 body + 143 hand / finger + 10 AUs) (Pavlakos et al., 2019) into a compact latent sub-space learned jointly with the decoder, preserving 99.3 % motion variance while enabling fast sampling. The MDN models
$p(z_{t}\mid z_{<t},H)=\sum_{k=1}^{K}\pi_{k}\,\mathcal{N}\bigl{(}z_{t}\mid\mu_{k},\sigma_{k}^{2}I\bigr{)},\quad K=5,$
capturing multimodal motion distributions (Bishop, 1994). Concurrently, the decoder outputs gloss logits $g_{t}$ (vocabulary $\sim 3$k, cross-entropy loss) and AU logits $a_{t}$ (7 classes, Focal loss) via multi-task heads (Lin et al., 2017). The joint training objective is
$\mathcal{L}=\lambda_{1}[\mathcal{L}_{\text{body}}+3\,\mathcal{L}_{\text{hand}}]+\lambda_{2}\mathcal{L}_{\text{Gloss}}+\lambda_{3}\mathcal{L}_{\text{AU}}$
where the suggested values of $\lambda$ are:
$(\lambda_{1},\lambda_{2},\lambda_{3})=(1,0.6,0.4),$
where hand-joint errors receive a 3×3$\times$3 × weight to emphasize clarity of critical articulations. The pipeline is shown in Figure 2.
Figure 2: Streaming Conformer–Transformer–MDN pipeline.
3.3 Editable Transformer Architecture in Sign Language Motion Generation
To bridge model inference and human agency, we introduce a structured JSON intermediate layer that explicitly exposes the core parameters of each sign unit—gloss label, start/end timestamps, handshape, motion trajectory, facial expression, and syntactic role—rather than treating generation as a “black box.” Crucially, this schema (gloss, start, end, handshape, movement, expression, syntax_tag, etc.) was not arbitrarily defined by developers but co‐created through two rounds of card sorting, priority voting, and participatory workshops with deaf users and professional interpreters. This ensures that the JSON fields align with users’ mental models and supports intuitive manipulation in the front‐end editor.
Figure 3: Resampling‐Hook partial re-sampling workflow.
In the live interface, each JSON entry is bound to the corresponding latent sequence segment via the “Resampling Hook” in Algorithm 2. Users modify any field using drag-and-drop, sliders, or dropdowns; the system detects the edit, injects the updated latent $\hat{z}_{t-1}$ into the autoregressive Transformer-MDN decoder, and efficiently recomputes only the affected subsequence $\{z_{t},\dots,z_{T}\}$. This partial resampling strategy preserves overall fluency while delivering sub-100 ms responsiveness.
  "gloss": "THANK-YOU",
  "start": 0.45,
  "end": 1.08,
  "handshape": "flat",
  "movement": "forward",
  "expression": "smile",
  "syntax_tag": "statement"
To further enhance explainability, we map the MDN mixture weights $\{\pi_{k}\}$ onto the 3D skeleton as a per‐frame heatmap: higher $\pi_{k}$ values render with greater opacity, enabling users to pinpoint segments where the model’s uncertainty or multimodality is greatest and focus their edits there. All UI components—including heatmap toggles and JSON fields—support keyboard, voice, and assistive‐device inputs in compliance with WCAG 2.2 AA, ensuring equitable access across diverse user abilities.
Algorithm 2 Interactive Transformer-MDN Decoder with Resampling Hook
  Input: features $\mathbf{H}$, previous latent $z_{t-1}$ or user‐edited $\hat{z}_{t-1}$
  Output: latent $z_{t}$, gloss $g_{t}$, AU $a_{t}$
  // 1. Self- and Cross-Attention
  $q_{t}\leftarrow\mathrm{SelfAttn}(z_{<t})$
  $c_{t}\leftarrow\mathrm{CrossAttn}(q_{t},\mathbf{H})$
  $s_{t}\leftarrow\mathrm{FFN}(q_{t}+c_{t})$
  // 2. MDN Prediction
  [$[\{\pi_{k},\mu_{k},\Sigma_{k}\}]\leftarrow\mathrm{MDNHead}(s_{t})$
  // 3. Sampling or Teacher‐Forcing
  if training then
     $z_{t}\leftarrow$ ground‐truth latent
  else
     $z_{t}\sim\sum_{k}\pi_{k}\mathcal{N}(\mu_{k},\Sigma_{k})$
  end if
  // 4. Gloss & AU
  $g_{t}\leftarrow\mathrm{GlossHead}(s_{t})$
  $a_{t}\leftarrow\mathrm{AUHead}(s_{t})$
  // 5. Resampling Hook (inference-time only; no gradients propagated)
  if user edits segment containing $t$ then
     Recompute $z_{t},\dots,z_{T}$ with updated $\hat{z}_{<t}$// forward-pass resampling only
  end if
3.4 Unity3D Animation Rendering and Client-Side Optimization
The generated motion key points are mapped and bound to the Humanoid Rig skeleton in Unity3D. We employ Two-Bone IK algorithms (Hecker et al., 2008) and Spline interpolation smoothing to further enhance motion naturalness and physical plausibility. On the inference side, the model utilizes 30% weight pruning, INT8 quantization, and TensorRT acceleration, reducing average frame time to $13ms$ (RTX 4070 Mobile). Together with (i) audio feature extraction $\approx 7ms$, (ii) Conformer encoding $\approx 30ms$, (iii) inverse kinematics $\approx 18ms$, and (iv) Unity rendering $\approx 35ms$, the end-to-end speech-to-avatar delay is $103\pm 6ms$, comfortably below our $150ms$ target. Even on standard notebook CPUs, it maintains stable performance at 15-25 FPS, enabling practical deployment on edge devices.
3.5 Human-in-the-Loop Optimization
To continuously align the model with real user needs, we embed a closed-loop feedback mechanism in production:
Feedback collection & model fine-tuning
After each generation or edit session, users rate the animation on a 5-point Likert scale, and all JSON diffs are logged. Weekly, professional interpreters annotate selected historic segments for terminology and grammatical accuracy. We assemble triplets $\bigl{(}\mathcal{J}^{\mathrm{orig}},\mathcal{J}^{\mathrm{edit}},r_{u},r_{e}\bigr{)}$—original JSON, user revision, user rating $r_{u}$, expert rating $r_{e}$—as incremental training data. The decoder parameters $\theta$ are then fine-tuned by minimizing a KL-regularized multi-task loss, combined with a PPO-style reward:
$J(\theta)=\mathbb{E}_{\pi_{\theta}}\!\Bigl{[}\sum_{t=0}^{\infty}\gamma^{t}\,R_{\phi}(s_{t},a_{t})\Bigr{]},R_{\phi}(s_{t},a_{t})=w_{u}\,r_{u}+w_{e}\,r_{e},$
where $\mathrm{D_{KL}}$-regularization encourages the updated policy $\pi_{\theta}$ to remain close to the pretrained one, and $(w_{u},w_{e})$ balance user versus expert signals (Schulman et al., 2017). Empirically, we perform micro-batches of fine-tuning every two weeks.
Design process & data governance
Our development follows the Double Diamond model (Design Council, 2005):
  • Discover & Define: user co-creation workshops, pain-point mapping, prioritization;
  • Develop & Deliver: prototype testing, expert review, data-driven iteration.
All audio/video samples undergo face-blurring and skeletal abstraction, then SHA-256 hashing for de-identification, in compliance with GDPR Art. 9(4) and CC-BY-NC 4.0.
4 Evaluation Methods and Results
This section systematically presents the multidimensional performance of our human-centered AI sign language animation system among real user groups. We combine quantitative and qualitative methods, focusing on key metrics such as usability, explainability, trustworthiness, editing burden, and inclusivity. The evaluation balances practical engineering metrics with design-science rigor.
4.1 Evaluation objectives, participants, and experimental procedures
Our evaluation objectives fall into two categories: (1) quantifying the system’s performance in terms of comprehension, naturalness, controllability, trustworthiness, and editing load; (2) qualitative analysis of how participatory information architecture and human-centered closed-loop optimization enhance user experience and actual sign language production workflows. The experimental procedure employs a Latin square balanced design to control order effects, with all tests conducted in quiet environments equipped with standard PCs and Unity animation preview interfaces.
Table 1: Experimental participants and grouping information
Group Number Age Range Gender Ratio Role Years of SL Use
(F/M) (Mean±$\pm$±SD)
Deaf Users 20 19–56 10/10 Users (n=20) $18.3\pm 7.2$
Interpreters 5 25–41 3/2 Professional Interpreters (n=5) $12.6\pm 4.5$
Total 25 19–56 13/12
Each participant sequentially completes two sets of tasks: ”Auto-generation (Auto)” and ”Generation + Editing (Edit)”, with each set containing 8 typical dialogue tasks (such as greetings, instructions, terminology, emotional expression, etc.), totalling 16 rounds of interaction. After each round, participants immediately fill out a Likert scale questionnaire and participate in a semi-structured interview, with the entire process being audio and video recorded. Both the scale and task order are controlled for sequence bias using a Latin square design.
4.2 Quantitative indicators and measurement tools
We comprehensively adopted multi-dimensional evaluation scales including system usability, cognitive load, trust and controllability, with quantitative analysis as follows:
  • Comprehensibility (C1–C4, Likert 1–5): Users’ subjective assessment of animation semantic accuracy;
  • Naturalness (C5–C8, Likert 1–5): Motion fluidity and facial expression naturalness;
  • System Usability (SUS, C9–C18, 0–100): Standard system usability score
  • Explainability & Controllability (C19–C26, Likert 1–5): Control capability over JSON structure and interaction flow;
  • Trust & Satisfaction (C27–C30, Likert 1–5): Trust in AI output results and overall satisfaction;
  • Cognitive Load (NASA-TLX simplified version, C31–C34, 0–100): Mental demand, physical demand, temporal demand, and overall burden
Additionally, we recorded completion time per task, edit counts, and distribution of frequently edited fields.
Table 2: Quantitative evaluation of main indicators (Auto vs. Edit)
Metric Mode Mean $\pm$ SD $t$ $p$ $d$ Change
Comprehensibility (C1–C4) Auto $3.2\pm 0.6$ 6.3 <.001 1.3 +28%
Edit $4.1\pm 0.5$
Naturalness (C5–C8) Auto $3.4\pm 0.5$ 5.8 <.001 1.2 +24%
Edit $4.2\pm 0.4$
System Usability (SUS) Auto $68.5\pm 8.1$ 5.1 <.001 1.0 +19%
Edit $81.3\pm 6.4$
Explainability & Controllability (C19–C26) Auto $3.1\pm 0.7$ 7.1 <.001 1.5 +32%
Edit $4.1\pm 0.6$
Trust & Satisfaction (C27–C30) Auto $2.9\pm 0.9$ 5.4 <.001 1.1 +34%
Edit $3.9\pm 0.7$
Cognitive Load (NASA-TLX) Auto $43.2\pm 11.0$ 2.9 0.008 0.6 –16%
Edit $36.5\pm 9.4$
Editing behavior analysis shows that users make an average of 1.7 edits per sentence in Edit mode, with the most frequent being hand gestures ($42\%$), duration ($28\%$), and facial expressions ($19\%$), while other fields (such as syntactic markers) account for relatively lower proportions. The average editing time per sentence is $7.8\pm 2.3$ seconds.
[…]
4.3 Explainability and cognitive transparency
To comprehensively evaluate the interpretability of AI systems and user understanding, we established three metrics: Explanation Satisfaction Score (ESS), Mental Model Accuracy (MMA), and Expected Calibration Error (ECE), which quantitatively reflect the system’s actual effectiveness in improving cognitive transparency. The specific results are shown in Table 3.
Table 3: Quantitative evaluation of explainability and transparency
Metric Mode Mean $\pm$ SD $t$ $p$ $d$ Change
ESS (1–5) Auto $3.1\pm 0.7$ 5.2 <.001 1.0 +29%
Edit $4.0\pm 0.6$
MMA (0–5) Auto $2.2\pm 1.1$ 4.5 <.001 0.9 +45%
Edit $3.2\pm 0.9$
ECE (%) Auto 12.4 4.2 <.001 0.85 –40%
Edit 7.5
The data shows that in Edit mode, the average ESS score increased from $3.1$ to $4.0$, with significant growth in MMA as well. This indicates that structured JSON and visual explanation mechanisms helped users clearly understand the system’s reasoning and action generation process, greatly reducing the black box feeling of AI. Meanwhile, ECE decreased from $12.4\%$ to $7.5\%$, meaning users’ confidence in the system’s output became more aligned with its actual performance, avoiding issues of over-trust or blind skepticism.
[…]
Overall, Edit mode improves interpretability and cognitive transparency by allowing users to not only understand the system but also effectively articulate its workings. This significantly strengthens users’ sense of control and trust foundation in AI systems—one of the key values pursued by human-centered AI.
4.4 Fairness, inclusivity, and green sustainability
This section focuses on analyzing the system’s performance in group fairness and energy efficiency optimization. Experimental results are shown in Table 4:
Table 4: Fairness and energy efficiency metrics (Auto vs. Edit)
Metric Auto Edit Improvement Rate
Demographic Gap (Gender) 0.42 0.18 57%
Demographic Gap (Age) 0.38 0.16 58%
Energy Efficiency E/F (J/frame) 0.24 0.17 29%
In Edit mode, the score differences in key experience metrics (such as usability, comprehension, etc.) between gender and age groups significantly decreased. The Demographic Gap (gender) dropped from $0.42$ to $0.18$, and the Demographic Gap (age) decreased from $0.38$ to $0.16$, both showing over $50\%$ improvement. These results indicate that structured intermediate layers and inclusive interface design can effectively reduce experience disparities among different user groups, enhancing overall system fairness. This trend was also confirmed in ANOVA analysis, with significantly reduced variance in subjective scores between groups in Edit mode ($p<0.05$).
Regarding energy efficiency, the average energy consumption per frame for on-device inference decreased from $0.24J$ to $0.17J$, a $29\%$ reduction. This improvement primarily stems from model pruning, quantization, and efficient inference optimization, enabling the system to maintain smooth interactions while better adapting to energy-constrained mobile or embedded scenarios. The data validates the synergistic advantages of human-centered AI design in improving both fairness and sustainability.
4.5 Human-machine co-creation experience and sense of autonomy
Table 5 shows key quantitative metrics of the system in human-AI co-creation experience and user autonomy (Cohen’s d for paired samples was computed as the mean difference divided by the pooled SD, with Hedges’ g correction applied for small-sample bias.):
Table 5: Quantitative evaluation of human–machine co‐creation metrics
Metric Mode Mean $\pm$ SD $t$ $p$ $d$ Change
Sense of Agency (SoA) Auto $2.9\pm 0.8$ 6.0 <.001 1.2 +38%
Edit $4.0\pm 0.6$
Error Recovery Time (s) Auto $5.4\pm 1.9$ 8.2 <.001 1.7 –46%
Edit $2.9\pm 1.2$
Learning Curve Slope $\beta$ $+0.11\!\uparrow$ 3.9 <.001
Co-Creation Gain (CCU %) All Tasks 19% +19%
[…]
Integrating these findings, Edit mode not only significantly enhanced user control and collaboration efficiency but also demonstrated outstanding performance in lowering onboarding barriers and operational burdens. All key metrics showed statistically significant improvements, validating the theoretical hypothesis of human-centered AI empowering co-creation and autonomy.
4.6 Emotional resonance and multimodal expression
This section focuses on the performance of AI systems in multi-channel (hand and non-hand) sign language generation, with particular emphasis on quantitative improvements in facial expressions and emotional representation. We employ AU consistency rate (agreement between automated actions and manually annotated action units) and subjective emotional resonance scores (Likert 1-5) as evaluation metrics, with results shown in Table 6.
Table 6: Evaluation of multimodal expressiveness metrics
Metric Auto Edit Improvement
AU Accuracy Rate 71% 81% +10%
Emotional Resonance (Likert 1–5 scale) 3.3 4.0 +21%
In Edit mode, the AU (Action Unit) consistency rate increased by 10 percentage points, indicating that the system’s generation of facial expressions and other non-hand markers aligns more closely with manual reference results. Additionally, the subjective emotional resonance score rose from $3.3$ to $4.0$, a $21\%$ improvement, demonstrating that users perceive animations in Edit mode as more accurately conveying emotions and tone, with significantly enhanced naturalness and warmth in communication.
Correlation analysis further revealed a moderate positive relationship between AU consistency and subjective emotional scores ($r=0.56$,$p<0.01$), suggesting that the accuracy of multimodal facial and body generation directly impacts users’ recognition of the system’s expressive richness. In expert annotations, Edit mode also showed marked superiority over Auto mode in the completeness and subtlety of non-hand signals when conveying emotional nuances and professional contexts. These results collectively demonstrate that improvements in multimodal generation not only enrich the expressiveness of AI sign language animations but also provide robust data-driven support for the inclusivity and effectiveness of accessible interaction systems.
In summary, Edit mode enhances the coordination of multi-channel generation and emotional expressiveness, enabling the system to better meet the dual demands of ”content + emotion” in real-world communication scenarios. This progress underscores the practical value of human-centered AI in multimodal interaction and social communication.
4.7 Qualitative evaluation and expert feedback
The qualitative section employed a multi-method approach combining semi-structured interviews, expert annotations, and open-text feedback to conduct thematic analysis (NVivo coding) of users’ authentic experiences and pain points. Key themes emerged:
  • T1. Control and Trust Enhancement: ”The editor lets me dictate animation details—the system feels more like an assistant.”
  • T2. Low Learning Curve: ”Five minutes to get started, intuitive interface, seamless edits.”
  • T3. Feature Refinement Requests: ”Need richer facial expressions; better predictive text for technical terms.”
  • T4. Emotional Journey: ”Satisfaction grows during the preview-edit-repreview cycle, but frustration arises when auto-mode makes errors.”
Expert annotation results are shown in Table 7.
Table 7: Expert annotation results
Metric Auto Edit Absolute Change
Term Accuracy 76% 88% +12%
Non-hand Signal Omission Rate 24% 8% –16%
Expert Consensus Score (1–5 scale) 3.5/5 4.4/5 +0.9
Expert annotation results, as summarized in Table 7, quantitatively confirm these qualitative insights: Edit mode demonstrated a $12\%$ increase in term accuracy, a $16\%$ reduction in non-hand signal omissions, and a 0.9-point gain in expert consensus score. These improvements reflect both the enhanced semantic precision and the expanded non-manual expressiveness achieved via structured editing and participatory feedback mechanisms.
Both users and experts emphasized the seamless integration between the JSON editor and animation preview, which facilitated rapid correction and iterative refinement. Future development should focus on further expanding the range of facial action units (AUs) and implementing intelligent auto-completion for specialized vocabulary, as highlighted in both open feedback and expert suggestions.
In summary, the qualitative and expert findings corroborate the quantitative trends reported in Sections 4.2–4.6, underscoring the value of human-in-the-loop, explainable, and co-creative system design for accessible and expressive sign language animation.
5 Discussion
5.1 Key Findings and Implications
Our experimental results robustly validate the practical value of a human-centered approach in speech-to-sign-language generation. The integration of a structured JSON intermediate representation and interactive editor yields significant gains in comprehension, naturalness, and usability (SUS), while also enhancing interpretability, controllability, and user trust. Edit mode empowers users to promptly correct errors, tailor outputs to personal linguistic habits, and maintain smooth communication—all with minimal additional cognitive load, as evidenced by NASA-TLX scores.
Statistical analysis further shows that controllability and interpretability are strong predictors of trust, highlighting the importance of user agency in AI-assisted communication. Qualitative feedback and expert annotations underscore these findings, confirming that participatory workflows not only reduce translation errors and omissions of non-manual information, but also foster inclusivity and professional reliability. The combined quantitative and qualitative evidence establishes a robust paradigm for future accessible, explainable, and user-adaptive sign language AI systems.
5.2 Limitations
Despite these advances, several limitations remain:
  • Nuanced Expression: Current models capture only core actions and primary facial expressions, with limited support for subtle emotions, spatial rhetoric, and personalized sign styles.
  • Non-Manual Coverage: Automated generation does not yet include full-body non-manual signals such as shoulder movement, body posture, and gaze, limiting expressiveness for complex semantics and grammar.
  • Editor Extensibility: The editor currently supports only basic fields; fine-grained editing for parameters such as intensity, orientation, and speed is not yet implemented.
  • Sample Diversity: User studies, while diverse in gender and age, remain limited in scale and regional coverage, with further work needed for international and dialectal adaptation.
  • Edge Device Adaptability: Latency and stability have not been fully validated on low-end devices and in poor network environments.
5.3 Future Directions
Future work should address these limitations by:
  • Diversifying Action Generation: Incorporate style transfer, emotion tagging, and diversity sampling to enable richer, more expressive sign animation for literature, performance, and multicultural contexts.
  • Advancing Multimodal Data and Annotation: Expand datasets to cover full-body, micro-expressions, and eye gaze; refine hierarchical annotation systems for better non-manual signal learning.
  • Intelligent and Personalized Editing: Develop adaptive editor features—such as auto-completion, grammar correction, and style archiving—for personalized, accessible, and inclusive interaction.
  • Real-World and Community Deployment: Collaborate with schools, organizations, and enterprises for long-term field studies, real-world deployment, and continuous learning.
  • Mobile Optimization: Research model compression, elastic architectures, and edge-cloud solutions to ensure low-latency and high reliability on mobile and resource-constrained devices.
5.4 Outlook
This research demonstrates both the theoretical value and real-world impact of human-centered AI in sign language generation. Ongoing efforts will focus on bridging design and technology, deepening collaboration with the deaf community and practitioners, and promoting the evolution of AI sign language tools from research prototypes to universally accessible infrastructure, advancing the cause of barrier-free communication and information equity.
Appendix A Evaluation Questionnaire Forms
A.1 Instructions
After each task (Auto or Edit), please rate your agreement with each statement on a 5‐point Likert scale: 1 = Strongly Disagree  2 = Disagree  3 = Neutral  4 = Agree  5 = Strongly Agree
A.2 Comprehensibility (C1–C4)
Item Statement
C1 The animation accurately conveyed the intended meaning of the input sentence.
C2 I could understand each sign’s meaning without additional explanation.
C3 The facial expressions and non-manual markers matched the semantic intent of the source message.
C4 Overall, I did not need to guess or infer extra context to grasp the message.
A.3 Naturalness (C5–C8)
Item Statement
C5 The hand movements appeared smooth and continuous.
C6 Transitions between consecutive signs felt natural.
C7 Facial expressions (e.g. eyebrow raises, mouth movements) looked realistic.
C8 The overall avatar motion seemed human-like rather than robotic.
A.4 Explainability & Controllability (C19–C26)
Item Statement
C19 I understood how edits in the JSON structure translated into changes in the animation.
C20 The intermediate representation (JSON) provided clear insight into the system’s decision process.
C21 I felt in control of the generation workflow at all times.
C22 I could easily manipulate parameters (e.g. handshape, trajectory) to customize the animation.
C23 The system’s feedback (visual preview) clearly indicated how my edits would affect the final animation.
C24 I was able to correct errors in the animation without confusion.
C25 The editing interface layout was intuitive for adjusting specific animation attributes.
C26 I felt confident that my changes would be accurately reflected when I replayed the animation.
A.5 Trust & Satisfaction (C27–C30)
Item Statement
C27 I trust this system to produce reliable and accurate sign language animations.
C28 I am satisfied with the overall quality of the generated animations.
C29 I would be comfortable using this system in my daily sign language production workflow.
C30 I feel confident sharing animations produced by this system with colleagues or clients.
A.6 System Usability Scale (SUS)
Please indicate your agreement (1–5) with the following:
1. I think that I would like to use this system frequently.
2. I found the system unnecessarily complex.
3. I thought the system was easy to use.
4. I think that I would need the support of a technical person to be able to use this system.
5. I found the various functions in this system were well integrated.
6. I thought there was too much inconsistency in this system.
7. I would imagine that most people would learn to use this system very quickly.
8. I found the system very cumbersome to use.
9. I felt very confident using the system.
10. I needed to learn a lot of things before I could get going with this system.
Scoring: Convert items 1,3,5,7,9 to 0–4 by subtracting 1; reverse‐score items 2,4,6,8,10; sum and multiply by 2.5 to yield 0–100.
A.7 NASA-TLX
The NASA-Task Load Index (TLX) assesses perceived workload on six dimensions using a two-step process: (1) weight derivation via pairwise comparisons; (2) workload rating. Compute overall workload as the weighted average of dimension ratings.
Step 1: Weight Derivation
Rank the relative importance of the six dimensions by indicating, for each pair, which dimension contributed more to your workload. Record your choices in the pairwise matrix below.
Mental Physical Temporal Performance Effort Frustration
Mental Demand
Physical Demand
Temporal Demand
Performance
Effort
Frustration
From this matrix, compute each dimension’s weight $W_{i}$ as the number of times it was chosen (range 0–5), then normalize:
$\tilde{W}_{i}=\frac{W_{i}}{\sum_{j=1}^{6}W_{j}},$
where $\sum\tilde{W}_{i}=1$.
Step 2: Workload Rating
Rate each dimension on a 0–100 scale:
Dimension Rating (0 = Low … 100 = High)
Mental Demand
Physical Demand
Temporal Demand
Performance
Effort
Frustration
Overall Workload Score
Compute the weighted workload:
$\mathrm{NASA\text{-}TLX}=\sum_{i=1}^{6}\tilde{W}_{i}\times R_{i},$
where $R_{i}$ is the 0–100 rating for dimension $i$.
This detailed procedure yields a scalar workload index reflecting both the importance and the perceived level of each demand dimension.“‘
Appendix B Interview Guide & Coding Scheme
B.1 Semi-Structured Interview Protocol
Participants were asked the following core questions after completing each condition. Probes (in italics) were used to elicit depth.
  1. 1.
    Overall Experience “How would you describe your overall experience with the system today?” Probe: Which part felt most intuitive or most challenging?
  2. 2.
    Control and Trust “Can you tell me about a moment when you felt in control of the animation?” Probe: Did any aspect make you doubt the system’s reliability?
  3. 3.
    Learning Curve “How quickly did you learn to perform edits?” Probe: Which features took longer to grasp, if any?
  4. 4.
    Error Handling “Describe how you fixed any mistakes made by the system.” Probe: How easy was it to identify and correct an error?
  5. 5.
    Emotional Response “How did the system’s animations affect your emotional engagement?” Probe: Did you feel more satisfied watching Edit mode vs. Auto mode?
  6. 6.
    Interface Feedback “What suggestions do you have for improving the editor or preview?” Probe: Are there any controls you wish were available?
B.2 NVivo Codebook
Code Definition Example Excerpt
Control & Trust Enhancement (T1) Instances where users describe feeling agency or confidence in system output “I always knew exactly what would happen when I changed the trajectory.”
Low Learning Curve (T2) References to ease and speed of initial adoption “I got the hang of the JSON editor in under five minutes.”
Feature Refinement Requests (T3) Suggestions for additional functionality or improvements “It would help to have predictive text for specialized signs.”
Emotional Journey (T4) Descriptions of emotional responses (e.g., satisfaction, frustration) “I felt frustrated when Auto mode made a wrong sign.”
B.3 Coding Procedure
  1. 1.
    Familiarization: Transcribe audio recordings verbatim and read through all transcripts.
  2. 2.
    Open Coding: Assign initial codes line-by-line in NVivo, allowing new themes to emerge.
  3. 3.
    Axial Coding: Group related open codes under the four pre-defined themes (T1–T4).
  4. 4.
    Selective Coding: Refine themes by merging or splitting codes to maximize internal homogeneity and external heterogeneity.
  5. 5.
    Inter-Rater Reliability: A second coder independently coded 20% of transcripts; Cohen’s $\kappa=0.82$.
Appendix C Experimental Design and Randomization
This appendix details the Latin‐square task ordering, counterbalancing scheme, and laboratory setup used to control for order and carryover effects.
C.1 Latin Square Task Ordering
We employed a 4×4 Latin square to balance the order of four task types across participants. Each row represents one of four participant groups.
| Group   | Task 1     | Task 2     | Task 3     | Task 4     |
| G1      | Greeting   | Instruction| Terminology| Emotion    |
| G2      | Instruction| Terminology| Emotion    | Greeting   |
| G3      | Terminology| Emotion    | Greeting   | Instruction|
| G4      | Emotion    | Greeting   | Instruction| Terminology|
Figure 4: Latin‐square assignment of the four dialogue tasks (Greeting, Instruction, Terminology, Emotion) across four groups (G1–G4).
C.2 Counterbalancing Scheme
Participants (N=25) were randomly assigned to one of the four Latin‐square groups. The following flowchart illustrates the randomization process:
        [ All Participants (N=25) ]
             Random Shuffle
        G1 (6)  G2 (6)  G3 (6)  G4 (7)
       Sequence1 Sequence2 Sequence3 Sequence4
Figure 5: Random assignment of participants into four groups (G1–G4) with approximately equal group sizes, each following a distinct task order (Sequences 1–4).
C.3 Experimental Environment Setup
All sessions were conducted in a quiet interview room. The participant sat at a desk facing a 24-inch monitor (1920×1080 px, 60 Hz) displaying the Unity 2023.3 animation preview. Directly beneath the monitor sat the experimental desktop (Windows 11, Intel i7-13700k CPU, 16 GB RAM, Nvidia RTX 4070), which ran both Unity and OBS Studio to capture synchronized screen, audio, and webcam video.
A Logitech C920 webcam (1080 p @ 30 fps) was mounted on a tripod 0.5 m above the top edge of the monitor, angled downward at 30° to capture the participant’s upper body and hands. All video and audio streams were recorded at 30 fps via OBS with lossless compression.
To prevent visual cues, a 30 cm high opaque divider was placed between the participant and the experimenter’s workstation. Ambient lighting was kept constant at 300 lux, and background noise was below 50 dB to ensure consistent recording quality.
Appendix D Energy Consumption and Performance Measurement
This appendix describes the measurement equipment, methods for synchronizing power and frame events, and the mobile/embedded deployment configurations used in our energy and performance evaluation.
D.1 Measurement Equipment and Methodology
  • Power Meter: Monsoon Power Monitor v3
    • Accuracy: ±$\pm 0.5\%$
    • Voltage range: 0–5 V DC
    • Sampling rate: 5 kHz (200 µs resolution)
    • Connection: inline to the device’s 5 V supply line
  • Logic Analyzer for Frame Sync: Saleae Logic Pro 16
    • Sample rate: 24 MHz
    • Channels:
        Channel 1: TTL “frame start” pulse generated by Unity via GPIO
        Channel 2: optional “inference start” marker
    • Used to align power trace with frame boundaries.
  • Data Capture Workflow:
    1. 1.
      Start Monsoon trace and Logic capture simultaneously.
    2. 2.
      Launch inference script; Unity emits a GPIO pulse at each frame presentation.
    3. 3.
      Stop capture after 1000 frames to ensure statistical significance.
    4. 4.
      Post‐process: parse TTL pulses to segment per‐frame energy $E_{i}$, compute average and standard deviation.
D.2 Mobile and Embedded Deployment Configurations
Budget and Platform Choices
All hardware was procured under a limited research budget ( \$200 USD per platform). We selected commodity devices with community support.
  • Smartphone (Mobile):
    • Model: Google Pixel 7 (Snapdragon 8 Gen 2)
    • OS: Android 13
    • Framework: TensorFlow Lite with NNAPI acceleration
    • Pruning: 30% filter‐level magnitude pruning applied in PyTorch prior to conversion
    • Quantization: Post‐training dynamic range quantization to INT8
    • Measurement: Monsoon inline at USB Type-C power, sampling at 5 kHz
  • Embedded (Edge):
    • Board: Raspberry Pi 4 Model B (4 GB RAM)
    • OS: Raspberry Pi OS (64-bit)
    • Framework: TensorFlow Lite with Edge TPU (Coral USB Accelerator)
    • Pruning: 25% structured channel pruning (TensorFlow Model Optimization Toolkit)
    • Quantization: Full integer quantization (weights + activations to INT8)
    • TPU Config: Edge TPU compiler v16.0, batch size = 1
    • Measurement: INA260 I²C power sensor (Adafruit breakout) at 2 kHz sampling, logged on Pi
D.3 Performance Metrics and Analysis
  • Per‐Frame Energy:
    $E_{\mathrm{frame}}=\frac{1}{N}\sum_{i=1}^{N}V_{i}\times I_{i}\times\Delta t,$
    where $V_{i}$, $I_{i}$ are instantaneous voltage/current samples during frame $i$, $\Delta t=200\,\mu\mathrm{s}$.
  • Inference Latency:
    • Measured from “inference start” TTL to “frame start” TTL
    • Reported as mean $\pm$ SD over 1,000 frames
  • CPU/GPU Utilization (Mobile):
    • Sampled via Android’s adb shell top at 100 ms intervals
    • Correlated with power trace to attribute energy to compute load