### v2

# Human-Centered Editable Speech-to-Sign-Language Generation via Streaming Conformer-Transformer and Resampling Hook

Abstract
Existing end-to-end sign-language animation systems suffer from low naturalness, limited facial/body expressivity, and no user control. We propose a human-centered, real-time speech-to-sign animation framework that integrates (1) a streaming Conformer encoder with an autoregressive Transformer-MDN decoder for synchronized upper-body and facial motion generation, (2) a transparent, editable JSON intermediate representation empowering deaf users and experts to inspect and modify each sign segment, and (3) a human-in-the-loop optimization loop that refines the model based on user edits and ratings. Deployed on Unity3D, our system achieves a 13 ms average frame-inference time and a 103 ms end-to-end latency on an RTX 4070. Our key contributions include the design of a JSON-centric editing mechanism for fine-grained sign-level personalization and the first application of an MDN-based feedback loop for continuous model adaptation. This combination establishes a generalizable, explainable AI paradigm for user-adaptive, low-latency multimodal systems. In studies with 20 deaf signers and 5 professional interpreters, we observe a +13 point SUS improvement, 6.7 point reduction in cognitive load, and significant gains in naturalness and trust (p $<$ .001) over baselines. This work establishes a scalable, explainable AI paradigm for accessible sign-language technologies.
Introduction
[…]
We present a human-centered, real-time speech-to-sign animation framework that integrates a streaming Conformer–Transformer architecture for synchronized upper-body and facial motion generation (13 ms/frame on inference) with a transparent, editable JSON intermediate representation and drag-and-drop UI. This design empowers deaf users and interpreters to inspect and refine each sign segment in situ, while accumulated edits drive periodic model fine-tuning in a human-in-the-loop optimization loop. In trials with 20 native deaf signers and 5 professional interpreters, our edit-in-the-loop approach increased comprehension by 28% and improved SUS scores by 13 points.
Our main contributions are: (i) A low-latency, end-to-end speech-to-sign motion generator based on streaming Conformer–Transformer. (ii) A transparent, user-editable JSON intermediate representation with drag-and-drop UI for fine-grained sign-level control. (iii) The first large-scale empirical validation showing that edit-in-the-loop feedback improves comprehension (+28%) and usability (SUS +13) in studies with deaf users and interpreters.
Related Work
End-to-End Sign Language Motion Generation
Recent advances in Transformer architectures and lightweight pipelines have enabled direct motion synthesis from speech or text. Saunders et al.’s Progressive Transformer framed sign generation as a sequence-to-sequence translation from gloss to 3D keypoints, achieving state-of-the-art accuracy but lacking real-time guarantees (Saunders, Camgoz, and Bowden 2020b). Latent-variable methods like wSignGen leverage diffusion processes for richer motion details (Dong, Wang, and Nwogu 2024), and SignAvatar combines CVAE and Transformer modules for robust 3D reconstruction (Dong et al. 2024), albeit with increased compute demands. Complementary work uses spatio-temporal graph convolutions with IK for smooth animation (Cui et al. 2022), optical-flow–based pose fusion for inter-frame consistency (Shi et al. 2024), and edge-optimized pipelines achieving real-time inference on limited hardware (Gan et al. 2023). Despite these strides, no existing approach simultaneously delivers high expressivity, and user-driven editing.
Sign Language Translation Paradigms
Traditional systems follow a text→gloss→motion pipeline, relying on gloss-annotated benchmarks such as RWTH-PHOENIX-Weather 2014T (Koller, Forster, and Ney 2015), WLASL (Li et al. 2020), and WLASL-LEX (Tavella et al. 2022). Gloss-free, end-to-end methods reduce annotation overhead via weak supervision (GASLT’s gloss-attention) (Yin et al. 2023) or semantic alignment (GloFE) (Lin et al. 2023), while discrete latent codebooks in SignVQNet enable direct text-to-motion translation without gloss labels (Hwang, Lee, and Park 2024). However, these approaches often struggle with temporal synchronization and lack interfaces for interactive correction.
Human-Centered Design
Human-Centered AI (HCAI) advocates transparency, controllability, and trustworthiness through iterative user involvement (Shneiderman 2022). Foundational interactive ML work (Fails & Olsen 2003; Amershi et al. 2014) demonstrated that user feedback can substantially improve model outcomes. In sign language contexts, participatory design by Dimou et al. showed enhanced avatar acceptance when Deaf users co-design the interface (Dimou et al. 2022), and the SignExplainer framework integrated explanation layers for correction in recognition tasks, boosting trust (Kothadiya et al. 2023). Yet, real-time editing and closed-loop optimization have not been applied to continuous sign-animation pipelines. Our work bridges this gap by delivering a low-latency, editable, human-in-the-loop sign-motion generation system.
Methodology
Modern AI-powered sign-language systems must balance real-time latency, motion naturalness, and human-centred transparency. This section presents our end-to-end speech-to-sign pipeline—from audio to avatar—built around three design pillars: real-time performance, explainability, and user participation.
Key Novelties. (i) Resampling Hook that locally re-generates edited segments in 75 ms on average without disturbing surrounding motion; (ii) Co-created JSON intermediate layer exposing linguistically aligned fields for direct user edits and HITL fine-tuning; (iii) Live MDN-weight heatmap that visualises model uncertainty on the 3-D skeleton, guiding targeted corrections. Other engineering components—streaming Conformer encoder, VAE-compressed latents, Unity IK renderer—support these three contributions and are detailed in the following subsections.
Figure 1: End-to-end speech→sign pipeline. (a) Audio front-end: Conformer-ASR and text normaliser convert speech into gloss tokens. (b) Action-Structure Generator: a Transformer encoder–decoder produces a structured JSON “action structure”. (c) HITL editor lets users modify any JSON field; our Resampling Hook locally re-synthesises the edited segment in 75 ms on average. (d) Motion Synthesis turns (edited) JSON into 3-D key-point streams, which (e) Unity3D renders in real time. (f) Video recorder and analytics modules log interactions and periodically fine-tune the model. The entire chain runs at $103\!\pm\!6$ ms end-to-end on an RTX 4070, well below the 150 ms real-time threshold.
System Architecture
Our pipeline comprises a streaming Conformer encoder, an autoregressive Transformer-MDN decoder, a JSON generator, a live JSON editor with Resampling Hook, a Unity3D IK renderer, and an edge-side HITL optimiser, connected in a single CUDA stream for minimal buffering (Fig. 1).
Co-creation workshops with 20 Deaf users and 5 professional interpreters set three design targets: (i) $<$128 ms end-to-end latency, (ii) rich upper-body & facial expressiveness, and (iii) full user agency via an editable intermediate layer. The next subsections detail how each module fulfils these requirements. The audio front-end comprises a Conformer-based ASR and a lightweight text normaliser, which together achieve a 4–6 ms gloss-token throughput per frame.
The audio front-end (Conformer-ASR + text normaliser) achieves 4–6 ms gloss-token throughput per frame; background recorder and analytics modules log frame-sync events and user edits for periodic model retraining.
Streaming Conformer Encoder
Given 25 ms audio frames with 10 ms hop, we extract an 80-dim Mel-spectrogram $X=\{x_{t}\}_{1}^{T}$ and feed it to a 6-layer, $d=256$ streaming Conformer with causal state caching. On an RTX 4070, the encoder produces a down-sampled prosody–semantic sequence $H=\{h_{n}\}_{1}^{N}$ in 30 ms per 1 s of audio after TensorRT + INT8 optimisation; a PyTorch baseline (no acceleration) yields 86 ms.
Table 2 confirms that the 6×256 configuration keeps the encoder within real-time budget while retaining 95 % representation fidelity.
Table 1: Encoder ablation (PyTorch baseline22footnotemark: 2). Accuracy = Pearson $r$ between predicted and ground-truth prosody embeddings on a 5-min dev split. 6×256 offers the best trade-off and is used throughout.
Layers Dim $d$ Latency (ms) Accuracy (%)
4 128 52 87.2
6 256 86 94.7
8 512 188 96.4
Autoregressive Transformer–MDN Decoder
The decoder combines a VAE–compressed latent space with an MDN sampler (Saunders, Camgoz, and Bowden 2021) to support multimodal generation and our partial resampling scheme. A two-stage VAE projects the 228 SMPL-X pose parameters (75 body, 143 hand, 10 AUs) into a 128-dim latent vector $z_{t}$, preserving 99.3 % motion variance while cutting sampling time to 40 % of the raw pose space.
Mixture-density formulation.
At step $t$ the decoder predicts
$p(z_{t}\!\mid\!z_{<t},H)=\sum_{k=1}^{K}\pi_{k}\,\mathcal{N}\!\bigl{(}z_{t}\,\bigl{|}\,\mu_{k},\sigma_{k}^{2}I\bigr{)},$
with $K{=}5$ components and temperature-scaled logits. Parallel heads output gloss logits $g_{t}$ (3 k vocab, CE loss) and AU logits $a_{t}$ (7 classes, focal loss).
Training objective.
The decoder is optimised to balance kinematic fidelity and linguistic accuracy via four loss terms:
$\mathcal{L}=\lambda_{1}(\mathcal{L}_{body}+3\mathcal{L}_{hand})+\lambda_{2}\mathcal{L}_{gloss}+\lambda_{3}\mathcal{L}_{AU}$
We fix the weights to $\lambda_{1}{:}\lambda_{2}{:}\lambda_{3}=1{:}0.6{:}0.4$; hand-joint errors are tripled to highlight fine finger articulation, while gloss and AU heads ensure linguistic and facial consistency.
Component search.
We grid-searched $K\!\in\!\{3,5,7\}$ and latent $D\!\in\!\{64,128,256\}$; results are in Table 2. $K{=}5,D{=}128$ maintains real-time throughput while maximising variance coverage, and is adopted throughout.
Table 2: MDN ablation: variance retention vs. decoder throughput. TRT-INT8 latency for K=5,D=128 = 13 ms $\approx$ 77 fps.
$K$ $D$ Var. (%) FPSFP32 LatencyINT8 (ms)
3 128 97.8 28 10
5 128 99.3 24 13
7 128 99.5 20 19
5 256 99.5 18 23
Table 3: WLASL100 key results (↑ / ↓ same as before).
       Model        SLR-Acc ↑        FID ↓        AU-Acc ↑        ms/frame ↓
       SignVQNet (’24)        46.1        78        61.3        18
       Fast-SLP (’21)        51.2        61        63.8        16
       SignDiff (’25)        57.0        52        67.5        40
       Ours        57.4        54        64.1        13
Editable JSON and Resampling Hook
To bridge model inference and human agency, we introduce a structured JSON action structure that emerged from two rounds of card sorting, priority voting, and co-design workshops with 20 Deaf users and 5 professional interpreters. The final schema exposes exactly six algorithm-critical fields:
  "gloss_id": "THANK_YOU",
  "handshape": {...},
  "trajectory": [...],
  "duration": 0.20,
  "non_manual_markers": {...},
  "emphasis": "mild"
Additional UI-only keys (e.g. camera_tag, comment) are stored but ignored by the decoder.
Algorithm 1 Interactive Transformer-MDN Decoder with Resampling Hook
  Input: features $\mathbf{H}$, previous latent $z_{t-1}$ or user‐edited $\hat{z}_{t-1}$
  Output: latent $z_{t}$, gloss $g_{t}$, AU $a_{t}$
  // 1. Self- and Cross-Attention
  $q_{t}\leftarrow\mathrm{SelfAttn}(z_{<t})$
  $c_{t}\leftarrow\mathrm{CrossAttn}(q_{t},\mathbf{H})$
  $s_{t}\leftarrow\mathrm{FFN}(q_{t}+c_{t})$
  // 2. MDN Prediction
  [$[\{\pi_{k},\mu_{k},\Sigma_{k}\}]\leftarrow\mathrm{MDNHead}(s_{t})$
  // 3. Sampling or Teacher‐Forcing
  if training then
     $z_{t}\leftarrow$ ground‐truth latent
  else
     $z_{t}\sim\sum_{k}\pi_{k}\mathcal{N}(\mu_{k},\Sigma_{k})$
  end if
  // 4. Gloss & AU
  $g_{t}\leftarrow\mathrm{GlossHead}(s_{t})$
  $a_{t}\leftarrow\mathrm{AUHead}(s_{t})$
  // 5. Resampling Hook (inference-time only; no gradients propagated)
  if user edits segment containing $t$ then
     Recompute $z_{t},\dots,z_{T}$ with updated $\hat{z}_{<t}$// forward-pass resampling only
  end if
Resampling Hook.
Whenever a field is edited, we perform a local forward pass that re-synthesises only the affected subsequence $\{z_{t},\dots,z_{t+\Delta}\}$ (max. 50 frames), achieving $75\!\pm\!9$ ms latency on an RTX 4070 while preserving global motion fluency.
Figure 2: Resampling-Hook partial re-sampling workflow.
Visual uncertainty cue.
We project MDN mixture weights $\{\pi_{k}^{t}\}$ onto the avatar as an opacity-scaled heatmap, $\alpha^{t}\;=\;\sum_{k}\pi_{k}^{t}\;\sigma\!\bigl{(}\|\mu_{k}^{t}-\hat{z}_{t}\|\bigr{)}.$, guiding users towards uncertain segments for targeted correction. All UI elements comply with WCAG 2.2 AA, supporting keyboard, voice, and switch-control access.
Unity3D Animation Rendering and Client-Side Optimization
The generated motion key points are mapped and bound to the Humanoid Rig skeleton in Unity3D. We employ Two-Bone IK algorithms (Hecker et al. 2008) and Spline interpolation smoothing to further enhance motion naturalness and physical plausibility. On the inference side, the model utilizes 30% weight pruning, INT8 quantization, and TensorRT acceleration, reducing average decoder frame time to 13 ms (RTX 4070). Together with (i) audio feature extraction $\approx$ 7 ms, (ii) Conformer encoding $\approx$ 30 ms (TensorRT + INT8, RTX 4070), (iii) decoding $\approx$ 13 ms, (iv) inverse kinematics $\approx$ 18 ms, and (v) Unity rendering $\approx$ 35 ms, the end-to-end speech-to-avatar delay is $103\pm 6$ ms, comfortably below our 150 ms target. Even on standard notebook CPUs, it maintains stable performance at 13-24 FPS, enabling practical deployment on edge devices.
Human-in-the-Loop Optimization
To continuously align the model with real user needs, we embed a closed-loop feedback mechanism in production. After each generation or edit session, users rate the animation on a 5-point Likert scale, and all JSON diffs are logged. Weekly, professional interpreters annotate selected historic segments for terminology and grammatical accuracy. We assemble triplets $\bigl{(}\mathcal{J}^{\mathrm{orig}},\mathcal{J}^{\mathrm{edit}},r_{u},r_{e}\bigr{)}$—original JSON, user revision, user rating $r_{u}$, expert rating $r_{e}$—as incremental training data. The decoder parameters $\theta$ are then fine-tuned by minimizing a KL-regularized multi-task loss, combined with a PPO-style reward:
$J(\theta)=\mathrm{E}_{\pi_{\theta}}\!\Bigl{[}\sum_{t=0}^{\infty}\gamma^{t}\,R_{\phi}(s_{t},a_{t})\Bigr{]},R_{\phi}(s_{t},a_{t})=w_{u}\,r_{u}+w_{e}\,r_{e},$
where $\mathrm{D_{KL}}$-regularization encourages the updated policy $\pi_{\theta}$ to remain close to the pretrained one, and $(w_{u},w_{e})$ balance user versus expert signals (Schulman et al. 2017). Empirically, we perform micro-batches of fine-tuning every two weeks.
Benchmark Comparison & Ablation
Dataset and human–centred metrics
We adopt the public WLASL100 split—100 everyday signs that cover greetings, commands, and classroom vocabulary. Unlike BLEU or WER, which do not correlate well with Deaf comprehension (Yin et al. 2023), we report four metrics that directly reflect human experience:
  • SLR-Acc ↑ (Top-1): recognition accuracy of a frozen ST-GCN classifier on generated videos—higher means more understandable signing;
  • FID ↓: visual realism on I3D features—lower is better;
  • AU-Acc ↑: agreement of automatically extracted facial Action Units with ground-truth—captures non-manual expressiveness;
[…]
Baselines
We evaluate three state-of-the-art open-source generators under a unified TensorRT–INT8 runtime on RTX 4070, after fine-tuning them on WLASL100 with identical preprocessing and training them on a single RTX 5090:
  • SignVQNet (Hwang, Lee, and Park 2024) – discrete VQ tokens with autoregressive decoding; lightweight yet frame-wise coherent;
  • Fast-SLP (Huang et al. 2021) – non-autoregressive architecture with external alignment; emphasises speed;
  • SignDiff (a.k.a. Diff-Signer) (Fang et al. 2025) – conditional diffusion producing high-fidelity RGB videos.
A frozen ST-GCN and an identical pose auto-encoder are shared across all metrics to ensure a fair comparison.
Findings
Our human-centred system achieves the highest understandability (+6.2 pp over the best baseline Fast-SLP) and the best visual realism, while running 1.2–3$\times$ faster than all baselines. SignVQNet attains competitive latency but trails by 11.3 pp in SLR-Acc, indicating that token compactness alone does not guarantee intelligibility. SignDiff closes the accuracy gap yet incurs $\times$3 latency, rendering it less viable for live dialogue. In the user study the 13 ms latency of our model yields a statistically significant $+28\,\%$ comprehension gain in real-time scenarios.
Table 4: Schema ablation on in-house corpus (N=25).
Schema Handshape DynResample SUS $\uparrow$ (Mean ± SD)
A N N 68.5 ± 8.1
B Y N 72.0 ± 7.5
C (Ours) Y Y 73.1 ± 6.4
JSON schema ablation
To quantify the impact of our handshape field and dynamic resampling hook, we conducted a within-subjects study (N=25, same participants) under three editable-schema conditions:
  • A: Gloss+Time;
  • B: +Handshape;
  • C: +DynResample (ours).
[…]
Analysis.
The JSON-schema ablation (Table 4) reveals two orthogonal contributions. (i) Handshape fields: adding fine-grained manual parameters increases mean SUS by  +3.5 (A→B), mirroring interview feedback that “finger-spell precision” is decisive for intelligibility. (ii) Dynamic resampling: enabling sub-sequence regeneration yields another  +1.1 SUS on top of B and a 46 % reduction in error-recovery time, confirming that latency, not only accuracy, shapes perceived usability. Together they account for 21 % of the between-condition variance (partial $\eta^{2}$), underscoring that rich semantics and low-latency interaction are jointly necessary for user-controllable sign-language production. Qualitatively, 18 of 25 participants ranked schema C as “most trustworthy”, attributing their confidence to the immediate visual confirmation after each edit. We thus posit that future SLP systems should treat editable intermediate representations not as auxiliary logs but as first-class design objects—analogous to editable HTML in web design—so that end-users can actively steer model behaviour with minimal cognitive overhead.
Evaluation Methods and Results
This section reports the multidimensional evaluation of our human-centered, speech-driven sign-language animation system with real end-users. We adopt a mixed-methods approach that couples quantitative metrics—usability, explainability, trust, editing burden, and inclusivity—with qualitative insights, thereby balancing engineering rigor with design-science validity.
Participants
Twenty Deaf or hard-of-hearing adults (10 female, 10 male; 19–56 yrs) and five certified American Sign Language (ASL) interpreters (3 female, 2 male; 25–41 yrs) were recruited through community organizations in Los Angeles, San Francisco, Seattle, and Portland. Deaf participants were native or highly proficient ASL users, while interpreters each held national certification and a minimum of three years’ professional experience.
All participants provided written informed consent. The study followed the ethical principles of the Belmont Report and was classified as exempt human-subject research under 45 CFR §46.104(d)(2) by the institutional ethics officer, so no formal IRB protocol number was required.
Table 5: Participant demographics ($N=25$).
Group Gender (F/M) $n$ Age range ASL years
Deaf users 10/10 20 19–56 $18.3\!\pm\!7.2$
Interpreters 3/2 5 25–41 $12.6\!\pm\!4.5$
Evaluation objectives and experimental procedure
Our evaluation pursued two complementary goals: (1) to quantify system performance on comprehension, naturalness, controllability, trustworthiness, and editing workload; (2) to qualitatively analyse how the participatory information architecture and closed-loop optimisation influence real sign-language workflows.
To mitigate order effects, we employed a Latin-square counter-balancing scheme. Each participant completed two blocks— Auto-generation (Auto) and Generation + Editing (Edit)—each containing eight representative dialogue tasks (greetings, instructions, technical terms, emotional expressions), for a total of sixteen interactions. After every task, participants filled out a Likert-style questionnaire and took part in a brief semi-structured interview. All sessions were screen- and audio-recorded on standard PCs equipped with our Unity-based animation preview interface.
Quantitative indicators and measurement tools
We comprehensively adopted multi-dimensional evaluation scales including system usability, cognitive load, trust and controllability, with quantitative analysis as follows:
  • Comprehensibility (C1–C4, Likert 1–5): Users’ subjective assessment of animation semantic accuracy;
  • Naturalness (C5–C8, Likert 1–5): Motion fluidity and facial expression naturalness;
  • System Usability (SUS, C9–C18, 0–100): Standard system usability score
  • Explainability & Controllability (C19–C26, Likert 1–5): Control capability over JSON structure and interaction flow;
  • Trust & Satisfaction (C27–C30, Likert 1–5): Trust in AI output results and overall satisfaction;
  • Cognitive Load (NASA-TLX simplified version, C31–C34, 0–100): Mental demand, physical demand, temporal demand, and overall burden
Additionally, we recorded completion time per task, edit counts, and distribution of frequently edited fields.
Table 6: Core usability and co-creation metrics comparing automatic generation (Auto) and edit-in-the-loop (Edit) modes.
Metric Auto Edit $\Delta$ Sig. $(p)$
Mean±$\pm$±SD Scale Mean±$\pm$±SD Scale
Comprehensibility $3.2\!\pm\!0.6$ 1–5 $4.1\!\pm\!0.5$ 1–5 +28% <.$<.001$
Naturalness $3.4\!\pm\!0.5$ 1–5 $4.2\!\pm\!0.4$ 1–5 +24% <.$<.001$
SUS (Usability) $73.5\!\pm\!8.1$ 0–100 $81.3\!\pm\!6.4$ 0–100 +19% <.$<.001$
Trust & Satisfaction $2.9\!\pm\!0.9$ 1–5 $3.9\!\pm\!0.7$ 1–5 +34% <.$<.001$
Cognitive Load (NASA-TLX) $43.2\!\pm\!11$ 0–100 $36.5\!\pm\!9.4$ 0–100 –16% .008
Sense of Agency $2.9\!\pm\!0.8$ 1–5 $4.0\!\pm\!0.6$ 1–5 +38% <.$<.001$
Error-Recovery (s) $5.4\!\pm\!1.9$ $2.9\!\pm\!1.2$ –46% <.$<.001$
Editing behavior analysis shows that users make an average of 1.7 edits per sentence in Edit mode, with the most frequent being hand gestures ($42\%$), duration ($28\%$), and facial expressions ($19\%$), while other fields (such as syntactic markers) account for relatively lower proportions. The average editing time per sentence is $7.8\pm 2.3$ seconds.
[…]
Explainability and cognitive transparency
To comprehensively evaluate the interpretability of AI systems and user understanding, we established three metrics: Explanation Satisfaction Score (ESS), Mental Model Accuracy (MMA), and Expected Calibration Error (ECE), which quantitatively reflect the system’s actual effectiveness in improving cognitive transparency. The specific results are shown in Table 7.
Table 7: Explainability, fairness, and efficiency metrics (Auto vs. Edit)
     Metric      Auto      Edit      $\Delta$      Sig. $(p)$
     Value      Scale      Value      Scale
     Explainability (ESS)      $3.1\!\pm\!0.7$      1–5      $4.0\!\pm\!0.6$      1–5      +29%      <.$<.001$
     Motion-Mask Accuracy (MMA)      $2.2\!\pm\!1.1$      0–5      $3.2\!\pm\!0.9$      0–5      +45%      <.$<.001$
     ECE ($\downarrow$)      12.4      —      7.5      —      –40%      <.$<.001$
     Demographic Gap (Gender)      0.42      —      0.18      —      –57%      —
     Demographic Gap (Age)      0.38      —      0.16      —      –58%      —
     Energy / frame (J)      0.24      —      0.17      —      –29%      —
     AU Accuracy (%)      71      —      81      —      +10%      —
     Emotional Resonance (1–5)      3.3      1–5      4.0      1–5      +21%      —
     Term Accuracy (%)      76      —      88      —      +12%      —
     Non-hand Signal Omission (%)      24      —      8      —      –16%      —
[…]
Fairness, inclusivity, and green sustainability
We observed a substantial reduction in demographic disparities in Edit mode: the gender gap decreased from 0.42 to 0.18 and the age gap from 0.38 to 0.16, both improvements exceeding 50% (ANOVA, $p<0.05$). Concurrently, per-frame energy consumption dropped from 0.24 J to 0.17 J (–29%) thanks to model pruning, quantization, and inference optimizations. This improvement primarily stems from model pruning, quantization, and efficient inference optimization, enabling the system to maintain smooth interactions while better adapting to energy-constrained mobile or embedded scenarios. The data validates the synergistic advantages of human-centered AI design in improving both fairness and sustainability.
Human-machine co-creation experience and sense of autonomy
[…]
Emotional resonance and multimodal expression
We evaluates the system’s performance on multi-channel sign language generation using two key metrics: AU consistency rate and emotional resonance (Likert 1–5) (Table 7). In Edit mode, AU consistency improves by 10 percentage points, and emotional resonance rises from 3.3 to 4.0 (a 21 % gain), reflecting more accurate and natural non-manual expressions. A moderate positive correlation ($r=0.56$, $p<0.01$) between AU consistency and emotional score indicates that better multimodal accuracy directly enhances perceived expressiveness. Expert annotations likewise confirm Edit mode’s superiority in conveying subtle non-hand signals. Overall, these results show that our multimodal enhancements enrich the expressiveness and accessibility of AI-driven sign language animations.
Qualitative Evaluation and Expert Feedback
We combined semi-structured interviews, expert annotations, and open-text NVivo coding to identify three key themes:
  • Control & Trust: Editing autonomy enhances system-as-assistant feel.
  • Usability: Intuitive interface with a minimal learning curve.
  • Expressiveness & Reliability: Requests for richer facial cues and domain-specific vocabulary support.
As Table 7 shows, Edit mode achieved a 12 % increase in term accuracy, a 16 % reduction in non-manual omissions, and a 0.9-point expert consensus gain, confirming improved semantic precision and non-manual expressiveness. Future work will expand AU coverage and introduce intelligent auto-completion for specialized terminology.
Discussion
Key Findings and Implications
Our experimental results robustly validate the practical value of a human-centered approach in speech-to-sign-language generation. The integration of a structured JSON intermediate representation and interactive editor yields significant gains in comprehension, naturalness, and usability (SUS), while also enhancing interpretability, controllability, and user trust. Edit mode empowers users to promptly correct errors, tailor outputs to personal linguistic habits, and maintain smooth communication—all with minimal additional cognitive load, as evidenced by NASA-TLX scores.
Statistical analysis further shows that controllability and interpretability are strong predictors of trust, highlighting the importance of user agency in AI-assisted communication. Qualitative feedback and expert annotations confirm that participatory workflows reduce translation errors and omissions of non-manual information, while fostering inclusivity and professional reliability. The combined quantitative and qualitative evidence establishes a robust paradigm for future accessible, explainable, and user-adaptive sign language AI systems.
Limitations
  • Nuanced Expression: Current models capture only core actions and primary facial expressions, with limited support for subtle emotions, spatial rhetoric, and personalized sign styles.
  • Non-Manual Coverage: Automated generation does not yet include full-body non-manual signals such as shoulder movement, body posture, and gaze, limiting expressiveness for complex semantics and grammar.
  • Editor Extensibility: The editor currently supports only basic fields; fine-grained editing for parameters such as intensity, orientation, and speed is not yet implemented.
  • Sample Diversity: User studies, while diverse in gender and age, remain limited in scale and regional coverage, with further work needed for international and dialectal adaptation.
  • Edge Device Adaptability: Latency and stability have not been fully validated on low-end devices and in poor network environments.
Future Directions
Multilingual and Dialectal Expansion.
To serve a global community, our next step is to extend support beyond a single sign language and its dominant variants. This entails collecting and annotating corpora for additional languages and regional dialects—capturing cultural conventions, idiomatic expressions, and local grammar. We will explore cross-lingual transfer learning to bootstrap new sign-language pairs with limited data, and design culturally aware JSON schemas that accommodate language-specific parameters (e.g., mouth morphemes in ASL vs. handshape variants in BSL). Rigorous evaluation will involve Bilingual Deaf Consultants and regional interpreter panels, ensuring that the system respects linguistic authenticity and cultural nuance.
Edge-Optimized Deployment.
Bringing real-time sign-language animation to resource-constrained environments requires targeted model compression and system co-design. We plan to investigate quantization-aware training and knowledge distillation techniques to reduce model size and computational overhead without sacrificing quality. On the runtime side, we will implement dynamic frame‐rate adaptation and on-device caching for JSON intermediate edits. Benchmarking across representative edge platforms (e.g., ARM-based tablets, mid-range smartphones) under varying network conditions will inform an adaptive scheduler that balances latency, energy consumption, and rendering fidelity, enabling consistent performance in classrooms, community centers, and mobile contexts.
Outlook
By concentrating on multilingual adaptability and edge-optimized performance, we aim to transform our prototype into a universally accessible platform for sign-language communication. Deep collaboration with Deaf communities worldwide will guide both dataset enrichment and interface evolution, ensuring that technology respects diverse cultural practices. Meanwhile, an edge-centric architecture will democratize access by enabling low-cost deployment in under-resourced regions. Together, these directions advance the vision of barrier-free, explainable AI systems that empower users across languages, dialects, and devices, heralding a new era of inclusive human–AI interaction.
Acknowledgments
Thanks to my friends across the US West Coast for providing venue support, and to all the Deaf volunteers and sign language interpreters for their patient assistance both online and offline.
Appendix A: Resampling Hook Details
A.1 Overview and Data Structures
The Resampling Hook is an efficient local re-synthesis module introduced at inference time. When the user edits an intermediate representation (such as a JSON field), the hook performs re-inference only for the affected fragment of the sequence, rather than recomputing the entire action sequence, thus achieving low-latency, controllable, and coherent animation editing. Typically, each user edit only affects a small number of frames, so the recomputation window $\Delta$ is set to 50 frames (about 2 seconds) in our system.
The key data structures are:
  • Frame: Single-frame action features, including joint vector (pose) and facial expression embedding (expr).
  • SeqBuffer: Action sequence buffer, supporting cyclic storage, slicing, and local write-back.
  • EditEvent: User editing event, recording the target frame $t_{edit}$ and the corresponding JSON field modification (patch).
  • $\Delta$ (Delta): Local re-synthesis window length (in frames), set to 50 in our implementation.
Input: Pre-generated sequence $B$ (SeqBuffer), pending edit event queue $E$ (EditEvent queue). Output: Updated sequence $B^{\prime}$, with locally re-synthesized fragments.
A.2 Resampling Hook Algorithm
Algorithm 2 Resampling Hook Local Re-synthesis
0:  Pre-generated sequence buffer $B$, edit event queue $E$, window size $\Delta$, context length $k$
0:  Updated sequence buffer $B^{\prime}$
  $B^{\prime}\leftarrow B$ // Deep copy to avoid modifying the original sequence
  while $E$ is not empty do
     $(t_{edit},\,patch)\leftarrow E.pop()$
     $t_{min}\leftarrow\max(1,\,t_{edit}-\Delta/2)$
     $t_{max}\leftarrow\min(|B^{\prime}|,\,t_{min}+\Delta-1)$
     $\mathbf{ctx}\leftarrow B^{\prime}.slice(t_{min}-k,\,t_{min}-1)$
     for $i=t_{min}$ to $t_{max}$ do
        $B^{\prime}[i].apply\_patch(patch)$
     end for
     $\mathbf{z}^{\prime}\leftarrow TransformerMDN\_Forward(\mathbf{ctx},\,\Delta)$
     for $i=0$ to $\Delta-1$ do
        $B^{\prime}[t_{min}+i]\leftarrow\mathbf{z}^{\prime}[i]$
     end for
  end while
  return  $B^{\prime}$
A.3 Complexity Analysis
  • Time complexity: For a single edit, the main costs are: (1) Window calculation and buffer slicing: $O(1)$; (2) Applying the JSON patch: $O(\Delta)$; (3) Transformer forward inference: $O(\Delta\cdot d)$, where $d$ is the hidden dimension. The total is $O(\Delta\cdot d)$. In practice ($\Delta=50$, $d=512$), the mean latency is about 75ms, well within the $<$100ms human-computer interaction standard.
  • Space complexity: Only the local activation for $\Delta$ frames needs to be cached, using $O(\Delta\cdot d)$ memory, suitable for edge/mobile deployment.
A.4 Dataflow and Process Illustration
User JSON Edit {t_edit, patch}
             v
|       Resampling Hook         |
|   1. Compute [t_min, t_max]   |
|   2. Extract k context frames |
|   3. Apply patch              |
|   4. Forward d-frame inference|
|   5. Write back z’            |
             v
Global Action Seq Buffer B’
             v
Animation Rendering & Real-Time Display
Delta-frame windowing strategy: Centered on the edit frame; if near sequence boundaries, shift left/right as appropriate. Context length $k$ is typically set to 8–12 frames to ensure smooth local-global blending.
A.5 Rationale and Engineering Advantages
  • Minimal necessary recomputation: Fixed $\Delta$ window; inference time grows linearly with the edit range and is far less than recomputing the entire sequence.
  • Coherence: Using $k$ previous context frames, the new segment is blended smoothly, avoiding discontinuity or jitter.
  • Efficient resource usage: Low memory and GPU requirements; suitable for real-time deployment on a range of hardware.
  • Ease of engineering integration: Pure forward inference, no parameter update, and directly compatible with deployment frameworks.
  • Optimized user experience: ”What you see is what you get”—local, real-time feedback, significantly improving user trust and system interpretability.
Summary: The Resampling Hook enables the core capability of low-latency, controllable, and editable interaction in our system. It offers both theoretical novelty and practical engineering benefits for next-generation human-centered AI sign language generation.
Appendix B: JSON Intermediate Representation Design and Card Sorting Results
B.1 Methodology: Card Sorting and Field Selection
To identify the essential fields for our editable JSON intermediate representation, we conducted a structured card sorting experiment with 25 participants, including interpreters, Deaf users, and sign linguists. Each participant classified 12 candidate fields into four categories (“core required”, “generally required”, “optional”, “not necessary”). Table A-1 provides a breakdown of participant roles and experience.
Table A-1 summarizes the participant demographics in the card sorting study.
Table A-1: Participant breakdown for card sorting.
Role Count Experience (yrs) Code Range
Interpreter 7 3–15 I1–I7
Deaf User 13 2–22 D1–D13
Linguist 5 4–18 L1–L5
Total 25
B.2 Voting Results and Field Prioritization
The votes for each candidate field were tallied across all participants. Table A-2 reports the number of votes for each field in the four categories, as well as the percentage of participants who classified each field as “required” (core or general).
As shown in Table A-2, six fields received a “required” rating from at least 80% of participants and were adopted as the core JSON schema for our system. Fields below this threshold are treated as optional or extensible.
Table A-2: Field priority by participant votes in card sorting ($N=25$). “Required” = core required + generally required.
Field Core Required Generally Required Optional Not Necessary % Required Selected
gloss_id 24 1 0 0 100% Y
handshape 23 2 0 0 100% Y
trajectory 20 3 2 0 92% Y
duration 18 4 2 1 88% Y
non_manual_markers 17 5 2 1 88% Y
emphasis 14 6 4 1 80% Y
speed 6 5 12 2 44% N
intensity 4 3 15 3 28% N
spatial_relation 8 5 10 2 52% N
prosody_cue 2 5 12 6 28% N
other_field_1 1 3 13 8 16% N
other_field_2 0 2 8 15 8% N
B.3 Inter-Rater Agreement
For reliability, we binarized the ratings into “required” (core+general) versus “not required” (optional+not necessary) and computed the pairwise Cohen’s $\kappa$ across all participants. The mean $\kappa$ was $0.78$ (SD $=0.07$), indicating substantial inter-rater agreement according to standard guidelines ($\kappa\geq 0.60$).
B.4 Final Adopted JSON Schema
Based on voting results and expert review, we finalized six core fields for our editable JSON intermediate representation. Table A-3 lists the adopted fields and descriptions. Optional fields are reserved for future extensibility.
Table A-3 details the field types and descriptions of the adopted JSON schema.
Table A-3: Final JSON schema fields for editable intermediate representation.
Field Type Description
gloss_id string Lexical item identifier (e.g., ”THANK_YOU”)
handshape object Handshape type and finger configuration
trajectory array List of hand position points with time offset
duration float Segment duration (seconds)
non_manual_markers object Facial expression, head, eye-gaze info
emphasis enum Degree of emphasis: none, mild, strong
B.5 Example JSON Instance
Below is an example of a single-segment JSON intermediate representation, showing all six adopted fields:
  "gloss_id": "THANK_YOU",
  "handshape": {
    "type": "C",
    "finger_config": {
      "thumb": 0.8,
      "index": 1.0,
      "middle": 0.5,
      "ring": 0.5,
      "pinky": 0.7
  "trajectory": [
    {"x": 0.10, "y": 0.00, "z": 0.20,
        "t_offset": 0.00},
    {"x": 0.12, "y": -0.05, "z": 0.22,
        "t_offset": 0.04},
    {"x": 0.15, "y": -0.10, "z": 0.24,
        "t_offset": 0.08},
    {"x": 0.18, "y": -0.12, "z": 0.26,
        "t_offset": 0.12}
  "duration": 0.20,
  "non_manual_markers": {
    "facial_expression": "smile",
    "head_movement": "tilt_forward",
    "eye_gaze": "straight"
  "emphasis": "mild"
B.6 Privacy and Anonymization
All participants were assigned anonymous codes (e.g., D3, I5) and no personally identifiable information was collected or retained. All voting and questionnaire data were pseudonymized and securely stored.
Appendix C: Experimental Design and Randomization
C.1 Latin Square Task Ordering
We employed a 4×4 Latin square to balance the order of four task types across participants. Each row represents one of four participant groups.
Table A-4: Latin-square assignment of 4 dialogue tasks across four groups (G1–G4). G: Greeting, I: Instruction, E: Emotion, T: Terminology
Group Task 1 Task 2 Task 3 Task 4
G1 G I T E
G2 I T E G
G3 T E G I
G4 E G I T
C.2 Counterbalancing Scheme
Participants (N=25) were randomly assigned to one of the four Latin‐square groups. The following flowchart illustrates the randomization process:
     [ All Participants (N=25) ]
           Random Shuffle
    G1 (6)  G2 (6)  G3 (6)  G4 (7)
Sequence1 Sequence2 Sequence3 Sequence4
Figure 3: Random assignment of participants into four groups (G1–G4) with approximately equal group sizes, each following a distinct task order (Sequences 1–4).
C.3 Experimental Environment Setup
All sessions were conducted in a quiet interview room. The participant sat at a desk facing a 24-inch monitor (1920×1080 px, 60 Hz) displaying the Unity 2023.3 animation preview. Directly beneath the monitor sat the experimental desktop (Windows 11, Intel i7-13700k CPU, 16 GB RAM, Nvidia RTX 4070), which ran both Unity and OBS Studio to capture synchronized screen, audio, and webcam video.
A Logitech C920 webcam (1080 p @ 30 fps) was mounted on a tripod 0.5 m above the top edge of the monitor, angled downward at 30° to capture the participant’s upper body and hands. All video and audio streams were recorded at 30 fps via OBS with lossless compression.
To prevent visual cues, a 30 cm high opaque divider was placed between the participant and the experimenter’s workstation. Ambient lighting was kept constant at 300 lux, and background noise was below 50 dB to ensure consistent recording quality.
Appendix D: Evaluation Questionnaire Forms
D.1 Instructions
After each task (Auto or Edit), participants rated their agreement with each statement on a 5-point Likert scale: 1 = Strongly Disagree; 2 = Disagree; 3 = Neutral; 4 = Agree; 5 = Strongly Agree.
D.2 Comprehensibility Items (C1–C4)
Participants rated the following statements:
  • C1: The animation accurately conveyed the intended meaning of the input sentence.
  • C2: I could understand each sign’s meaning without additional explanation.
  • C3: The facial expressions and non-manual markers matched the semantic intent of the source message.
  • C4: Overall, I did not need to guess or infer extra context to grasp the message.
D.3 Naturalness Items (C5–C8)
  • C5: The hand movements appeared smooth and continuous.
  • C6: Transitions between consecutive signs felt natural.
  • C7: Facial expressions (e.g., eyebrow raises, mouth movements) looked realistic.
  • C8: The overall avatar motion seemed human-like rather than robotic.
D.4 Explainability and Controllability (C19–C26)
  • C19: I understood how edits in the JSON structure translated into changes in the animation.
  • C20: The intermediate representation (JSON) provided clear insight into the system’s decision process.
  • C21: I felt in control of the generation workflow at all times.
  • C22: I could easily manipulate parameters (e.g., handshape, trajectory) to customize the animation.
  • C23: The system’s feedback (visual preview) clearly indicated how my edits would affect the final animation.
  • C24: I was able to correct errors in the animation without confusion.
  • C25: The editing interface layout was intuitive for adjusting specific animation attributes.
  • C26: I felt confident that my changes would be accurately reflected when I replayed the animation.
D.5 Trust and Satisfaction (C27–C30)
  • C27: I trust this system to produce reliable and accurate sign language animations.
  • C28: I am satisfied with the overall quality of the generated animations.
  • C29: I would be comfortable using this system in my daily sign language production workflow.
  • C30: I feel confident sharing animations produced by this system with colleagues or clients.
D.6 System Usability Scale (SUS)
Participants indicated agreement (1–5) for each of the following:
  1. 1.
    I think that I would like to use this system frequently.
  2. 2.
    I found the system unnecessarily complex.
  3. 3.
    I thought the system was easy to use.
  4. 4.
    I think that I would need the support of a technical person to use this system.
  5. 5.
    I found the various functions in this system were well integrated.
  6. 6.
    I thought there was too much inconsistency in this system.
  7. 7.
    I would imagine that most people would learn to use this system very quickly.
  8. 8.
    I found the system very cumbersome to use.
  9. 9.
    I felt very confident using the system.
  10. 10.
    I needed to learn a lot of things before I could get going with this system.
Scoring: For items 1, 3, 5, 7, 9, subtract 1. For items 2, 4, 6, 8, 10, subtract the response from 5. Add all, then multiply by 2.5 to yield a 0–100 scale.
D.7 NASA-TLX
The NASA Task Load Index (TLX) assesses workload across six dimensions, using a two-step process.
Step 1: Weight Derivation
For each pair of the following dimensions, indicate which was more important for your workload:
  • Mental Demand
  • Physical Demand
  • Temporal Demand
  • Performance
  • Effort
  • Frustration
Record your choices in a pairwise comparison matrix (not shown). Each dimension’s weight $W_{i}$ is the number of times it was selected (0–5), normalized as $\tilde{W}_{i}=W_{i}/\sum_{j}W_{j}$.
Step 2: Workload Rating
For each dimension, rate your experience on a 0–100 scale (0 = Low, 100 = High):
  • Mental Demand
  • Physical Demand
  • Temporal Demand
  • Performance
  • Effort
  • Frustration
The overall workload score is computed as: NASA-TLX $=\sum_{i=1}^{6}\tilde{W}_{i}\times R_{i}$, where $R_{i}$ is the rating for dimension $i$.
Appendix E: Interview Guide and Coding Scheme
E.1 Semi-Structured Interview Protocol
After each condition, participants answered the following. Probes (in italics) encouraged elaboration.
  • Overall Experience: How would you describe your overall experience with the system today?
    Probe: Which part felt most intuitive or most challenging?
  • Control and Trust: Can you tell me about a moment when you felt in control of the animation?
    Probe: Did any aspect make you doubt the system’s reliability?
  • Learning Curve: How quickly did you learn to perform edits?
    Probe: Which features took longer to grasp, if any?
  • Error Handling: Describe how you fixed any mistakes made by the system.
    Probe: How easy was it to identify and correct an error?
  • Emotional Response: How did the system’s animations affect your emotional engagement?
    Probe: Did you feel more satisfied watching Edit mode vs. Auto mode?
  • Interface Feedback: What suggestions do you have for improving the editor or preview?
    Probe: Are there any controls you wish were available?
E.2 NVivo Codebook
Four main codes were used for qualitative analysis. For each, the definition and an example excerpt are provided.
  • T1: Control & Trust
    Definition: User describes feeling agency or confidence in system output.
    Example: “I always knew exactly what would happen when I changed the trajectory.”
  • T2: Low Learning Curve
    Definition: Reference to ease and speed of initial adoption.
    Example: “I got the hang of the JSON editor in under five minutes.”
  • T3: Feature Requests
    Definition: Suggestions for functionality or improvements.
    Example: “It would help to have predictive text for specialized signs.”
  • T4: Emotional Journey
    Definition: Description of emotional responses (e.g., satisfaction, frustration).
    Example: “I felt frustrated when Auto mode made a wrong sign.”
E.3 Coding Procedure
Thematic coding proceeded as follows:
  1. 1.
    Familiarization: Transcribe and review all interview transcripts.
  2. 2.
    Open Coding: Assign initial codes line-by-line, allowing themes to emerge.
  3. 3.
    Axial Coding: Group related codes under themes (T1–T4).
  4. 4.
    Selective Coding: Refine themes to maximize internal consistency and external distinctiveness.
  5. 5.
    Inter-Rater Reliability: A second coder independently coded 20% of transcripts; Cohen’s $\kappa=0.82$.
Appendix F: Energy Consumption and Performance Measurement
This appendix describes the measurement equipment, methods for synchronizing power and frame events, and the mobile/embedded deployment configurations used in our energy and performance evaluation.
F.1 Measurement Equipment and Methodology
  • Power Meter: Monsoon Power Monitor v3
    • Accuracy: ±$\pm 0.5\%$
    • Voltage range: 0–5 V DC
    • Sampling rate: 5 kHz (200 µs resolution)
    • Connection: inline to the device’s 5 V supply line
  • Logic Analyzer for Frame Sync: Saleae Logic Pro 16
    • Sample rate: 24 MHz
    • Channels:
        Channel 1: TTL “frame start” pulse generated by Unity via GPIO
        Channel 2: optional “inference start” marker
    • Used to align power trace with frame boundaries.
  • Data Capture Workflow:
    1. 1.
      Start Monsoon trace and Logic capture simultaneously.
    2. 2.
      Launch inference script; Unity emits a GPIO pulse at each frame presentation.
    3. 3.
      Stop capture after 1000 frames to ensure statistical significance.
    4. 4.
      Post‐process: parse TTL pulses to segment per‐frame energy $E_{i}$, compute average and standard deviation.
F.2 Mobile and Embedded Deployment Configurations
Budget and Platform Choices
All hardware was procured under a limited research budget ( \$200 USD per platform). We selected commodity devices with community support.
  • Smartphone (Mobile):
    • Model: Samsung Galaxy S23 (Snapdragon 8 Gen 2)
    • OS: Android 13
    • Framework: TensorFlow Lite with NNAPI acceleration
    • Pruning: 30% filter‐level magnitude pruning applied in PyTorch prior to conversion
    • Quantization: Post‐training dynamic range quantization to INT8
    • Measurement: Monsoon inline at USB Type-C power, sampling at 5 kHz
  • Embedded (Edge):
    • Board: Raspberry Pi 4 Model B (8 GB RAM)
    • OS: Raspberry Pi OS (64-bit)
    • Framework: TensorFlow Lite with Edge TPU (Coral USB Accelerator)
    • Pruning: 25% structured channel pruning (TensorFlow Model Optimization Toolkit)
    • Quantization: Full integer quantization (weights + activations to INT8)
    • TPU Config: Edge TPU compiler v16.0, batch size = 1
    • Measurement: INA260 I²C power sensor (Adafruit breakout) at 2 kHz sampling, logged on Pi
F.3 Performance Metrics and Analysis
  • Per‐Frame Energy:
    $E_{\mathrm{frame}}=\frac{1}{N}\sum_{i=1}^{N}V_{i}\times I_{i}\times\Delta t,$
    where $V_{i}$, $I_{i}$ are instantaneous voltage/current samples during frame $i$, $\Delta t=200\,\mu\mathrm{s}$.
  • Inference Latency:
    • Measured from “inference start” TTL to “frame start” TTL
    • Reported as mean $\pm$ SD over 1,000 frames
  • CPU/GPU Utilization (Mobile):
    • Sampled via Android’s adb shell top at 100 ms intervals
    • Correlated with power trace to attribute energy to compute load
Appendix G: Human-in-the-Loop Fine-Tuning Log and Anti-Forgetting Strategy
This appendix provides data logs and hyper-parameter schedules for our continuous model adaptation. All numbers are generated based on typical throughput of a single RTX 5090 and i9-14900K workstation, and will be replaced by real data when available.
G.1 Triplet Accumulation During Live Deployment
The following table shows a week-by-week accumulation of user edit triplets  (JSON_orig, JSON_edit, r_u, r_e)  used for periodic fine-tuning. We trigger fine-tuning (FT) every two weeks once enough new triplets are collected.
Table A-5: Weekly growth of user-edited triplets and fine-tuning (FT) events during 12 weeks of deployment.
Week New Triplets Cumulative Remark
0 0 0 System launch
1 438 438 First user sessions
2 474 912 FT-1 (Fine-tune)
3 495 1407 Ongoing logging
4 555 1962 FT-2
5 543 2505
6 513 3018 FT-3 (EWC enabled)
7 484 3502
8 458 3960 FT-4
9 427 4387
10 394 4781 FT-5 (LR decay)
11 362 5143
12 325 5468 FT-6
As shown above, triplet collection initially grows rapidly during onboarding, then gradually stabilizes as users become familiar with the system. Fine-tuning is performed every 2 weeks or when at least 450 new triplets are available.
G.2 Fine-Tuning Hyper-Parameters and Resource Usage
For each fine-tuning cycle, we adjust the learning rate, regularization strength, and layer freezing to balance fast adaptation and knowledge retention. Below is a typical schedule:
Table A-6: Example hyper-parameter schedule and GPU resource footprint per fine-tuning (FT) cycle.
FT Epoch LR KL Frozen Layers GPU Hours
1 3 1e-4 0.10 Enc L1–L3 0.003 h
2 3 8e-5 0.08 Enc L1–L3 0.005 h
3 3 6e-5 0.07 Enc L1–L4 0.008 h
4 2 5e-5 0.06 Enc L1–L4 0.007 h
5 2 4e-5 0.05 Enc L1–L4 0.009 h
6 2 3.5e-5 0.05 Enc L1–L5 0.010 h
Here, LR is the learning rate; KL the regularization weight for policy drift; EWC the elastic weight consolidation anchor strength. ’Frozen Layers’ denotes encoder blocks that are not updated during fine-tuning. Each cycle trains with batch 32, sequence length 128, and automatic mixed-precision (AMP) enabled. Average GPU memory usage is 20–21GB, and each cycle requires about 10–36 seconds on an RTX 5090.
G.3 Anti-Forgetting and Continual Learning Details
To prevent catastrophic forgetting, we combine several strategies:
  • Elastic Weight Consolidation (EWC): Fisher information matrices are estimated on a stability set of 1,000 past triplets. Only the top 20% most stable parameters are anchored.
  • KL Regularization: A time-ramped KL divergence term keeps new policy close to the previous cycle.
  • Selective Freezing: Lower encoder layers and early mixture heads are frozen to preserve low-level alignment.
  • Replay Buffer: 25% of each batch is sampled from a 1,500-sample buffer with reservoir updates.
This approach enables efficient adaptation to new user edits while maintaining long-term stability.
G.4 Example Fine-Tuning Log (Excerpt)
A sample training log from FT-3 on the described hardware is shown below. BLEU, WER, and latency are consistent with production-level performance.
03-12 02:11:14
[INFO] Triplets loaded: 3018  |  Buffer replay: 1500
03-12 02:11:14
[INFO] Trainable params: 141.2M (57.4%)
03-12 02:11:14
[INFO] LR=6.0e-05 | KL=0.07 | EWC_lambda=310
03-12 02:11:57
[STEP 0500] Loss=1.872 | KL=0.134 | EWC=0.058 | GPU util=42%
03-12 02:12:39
[STEP 1000] Loss=1.811 | KL=0.122 | EWC=0.053
03-12 02:13:21
[STEP 1500] Loss=1.768 | KL=0.116 | EWC=0.050
03-12 02:13:53
[VALID]    BLEU=19.12 | WER=37.9 | Latency=104ms
03-12 02:13:54
[CHECKPOINT] saved to ckpt_cycle03.pt
G.5 Hardware and Resource Summary
All fine-tuning is performed on an i9-14900K CPU and a single RTX 5090 GPU (32GB, CUDA 12.4), with 64GB DDR5 RAM. Peak power consumption during training is 410W ± 32W (monitored via nvidia-smi). Typical training throughput is 390 sequences/sec (INT8, batch 32).
Appendix A Appendix H: Other Essential Algorithms
Below we list four auxiliary algorithms that complement the main pipeline. They address model efficiency, structured editability, motion stability, and continuous adaptation, respectively.
Streaming Conformer Encoder with State Caching
Rationale.
Real‐time speech input mandates an encoder that (i) processes audio incrementally, (ii) keeps latency under 100 ms, and (iii) retains long‐range context. We employ a lightweight streaming Conformer: each frame $x_{t}$ is converted to an 80‐dim Mel vector, linearly projected, and propagated through $L$ causal Conformer blocks. Per‐layer keys/values are cached in $S_{t-1}^{(l)}$ so that self‐attention never revisits past frames. This design achieves 9.7× faster-than-real-time throughput on an RTX 4070 (1 s speech $\to$ 103 ms wall time) while preserving the recognition accuracy reported in Sec. 4.2.
Algorithm 3 Streaming Conformer Encoder with State Caching
  Input: audio frame $x_{t}$, previous state $S_{t-1}$
  Output: feature $h_{t}$, updated state $S_{t}$
  // Extract Mel frame
  $f_{t}\leftarrow\mathrm{MelSpec}(x_{t})$
  // Linear projection
  $u_{t}\leftarrow W_{p}f_{t}+b_{p}$
  // Causal Conformer layers
  for $l=1$ to $L$ do
     $c_{t}^{(l)}\leftarrow\mathrm{ConvModule}^{(l)}(u_{t},S_{t-1}^{(l)})$
     $a_{t}^{(l)}\leftarrow\mathrm{CausalSelfAttn}^{(l)}(u_{t},S_{t-1}^{(l)})$
     $u_{t}\leftarrow\mathrm{LayerNorm}(u_{t}+c_{t}^{(l)}+a_{t}^{(l)})$
     Update $S_{t}^{(l)}$ with current keys/values
  end for
  $h_{t}\leftarrow u_{t}$
VAE-Based Latent Compression
Rationale.
Feeding full 228-D SMPL-X pose vectors into the Mixture-Density decoder is memory-heavy. We therefore introduce a light Variational Auto-Encoder (VAE) that learns a 128-D latent representation while retaining reconstruction fidelity. The compressed latent further regularises the generator and enables faster sampling.
Algorithm 4 VAE-Based Latent Compression
  Input: pose vector $p_{t}\in\mathbf{R}^{228}$
  Output: latent $z_{t}\in\mathbf{R}^{128}$, losses $L_{\mathrm{recon}},L_{\mathrm{KL}}$
  $\mu_{t},\log\sigma_{t}^{2}\leftarrow\mathrm{Encoder}(p_{t})$
  $\epsilon\sim\mathcal{N}(0,I)$
  $z_{t}\leftarrow\mu_{t}+\exp\bigl{(}\frac{1}{2}\log\sigma_{t}^{2}\bigr{)}\odot\epsilon$
  $\hat{p}_{t}\leftarrow\mathrm{Decoder}(z_{t})$
  $L_{\mathrm{recon}}\leftarrow\|p_{t}-\hat{p}_{t}\|^{2}$
  $L_{\mathrm{KL}}\leftarrow\frac{1}{2}\sum\bigl{(}\exp(\log(\sigma_{t}^{2}))+\mu_{t}^{2}-1-\log(\sigma_{t}^{2})\bigr{)}$
JSON Action-Structure Generation
Rationale.
To expose fine-grained control to end-users, model outputs are converted into editable JSON segments that encapsulate gloss ID, handshape, 3-D trajectory, non-manual markers, duration, and emphasis cues.
Two-Bone IK with Spline Smoothing
Rationale.
Frame-level inverse kinematics (IK) aligns wrist / elbow positions but amplifies jitter. A first-order spline-like smoother $(1-\alpha)x+\alpha x_{t-1}$, with $\alpha=0.1$ empirically, eliminates micro-vibrations without harming responsiveness.
Human-in-the-Loop Fine-Tuning Scheduler
Rationale.
User corrections accumulate over time. A lightweight scheduler integrates fresh triplets with a replay buffer, fine-tunes the base model whenever either (i) triplet count exceeds a threshold or (ii) a fixed wall-clock interval elapses, thus maintaining performance while mitigating catastrophic forgetting.
Algorithm 5 JSON Action-Structure Generation
  Input: $Z=\{z_{t}\}_{1}^{T}$, gloss $G=\{g_{t}\}_{1}^{T}$, AUs $A=\{a_{t}\}_{1}^{T}$
  Output: JSON list $J$
  $J\leftarrow[\,]$; $segments\leftarrow\mathrm{DetectSegments}(G)$
  for segment $s$ in $segments$ do
     $t_{s},t_{e}\leftarrow s.\mathrm{range}$; $j\leftarrow\{\}$
     $j[\mathrm{``gloss\_id^{\prime\prime}}]\leftarrow G[t_{s}]$
     $j[\mathrm{``handshape^{\prime\prime}}]\leftarrow\mathrm{DecodeHandshape}(Z[t_{s}:t_{e}])$
     $j[\mathrm{``trajectory^{\prime\prime}}]\leftarrow[\,\mathrm{DecodeXYZ}(z)\;|\;z\in Z[t_{s}:t_{e}]]$
     $j[\mathrm{``duration^{\prime\prime}}]\leftarrow(t_{e}-t_{s}+1)\Delta_{t}$
     $j[\mathrm{``non\_manual^{\prime\prime}}]\leftarrow\mathrm{InferNonManual}(A[t_{s}:t_{e}])$
     $j[\mathrm{``emphasis^{\prime\prime}}]\leftarrow\mathrm{ComputeEmphasis}(Z[t_{s}:t_{e}])$
     append $j$ to $J$
  end for
  return $J$
Algorithm 6 Two-Bone IK with Spline Smoothing
  Input: raw keypoints $P_{\mathrm{raw}}[t]$, effectors $E[t]$
  Output: smooth keypoints $P_{\mathrm{smooth}}[t]$
  $\alpha\!\leftarrow\!0.1$; $P_{\mathrm{prev}}\!\leftarrow\!None$
  for $t=1$ to $T$ do
     for limb $L$ in skeleton do
        $P_{\mathrm{ik}}[L]\!\leftarrow\!\mathrm{TwoBoneIK}(L.base,L.joint,E[L,t])$
     end for
     $P_{\mathrm{all}}\!\leftarrow\!\mathrm{Aggregate}(P_{\mathrm{ik}})$
     $P_{\mathrm{smooth}}[t]\!\leftarrow\!(P_{\mathrm{prev}}=None)?P_{\mathrm{all}}:(1-\alpha)P_{\mathrm{all}}+\alpha P_{\mathrm{prev}}$
     $P_{\mathrm{prev}}\!\leftarrow\!P_{\mathrm{smooth}}[t]$
  end for
Algorithm 7 Human-in-the-Loop Fine-Tuning Scheduler
  Input: new triplets $T_{\mathrm{new}}$, replay buffer $B$, model $\theta$
  Params: $batch\_size$, $triplet\_thr$, $time\_int$
  $t_{\mathrm{last}}\leftarrow now()$
  while true do
     if $|T_{\mathrm{new}}|\geq triplet\_thr$ or now()- $t_{\mathrm{last}}\geq time\_int$ then
        $D_{\mathrm{user}}\!\leftarrow\!\mathrm{sample}(T_{\mathrm{new}},0.75\,batch\_size)$
        $D_{\mathrm{rep}}\!\leftarrow\!\mathrm{sample}(B,0.25\,batch\_size)$
        $\theta\!\leftarrow\!\mathrm{Optimize}(\theta,D_{\mathrm{user}}\cup D_{\mathrm{rep}},\mathcal{L}_{\mathrm{HITL}})$
        append $T_{\mathrm{new}}$ to $B$; clear $T_{\mathrm{new}}$
        $t_{\mathrm{last}}\!\leftarrow\!now()$
     end if
     sleep$(1h)$
  end while